### v2

# An Efficient Learning Control Framework With Sim-to-Real for String-Type Artificial Muscle-Driven Robotic Systems

Abstract
Robotic systems driven by artificial muscles present unique challenges due to the nonlinear dynamics of actuators and the complex designs of mechanical structures. Traditional model-based controllers often struggle to achieve desired control performance in such systems. Deep reinforcement learning (DRL), a trending machine learning technique widely adopted in robot control, offers a promising alternative. However, integrating DRL into these robotic systems faces significant challenges, including the requirement for large amounts of training data and the inevitable sim-to-real gap when deployed to real-world robots. This paper proposes an efficient reinforcement learning control framework with sim-to-real transfer to address these challenges. Bootstrap and augmentation enhancements are designed to improve the data efficiency of baseline DRL algorithms, while a sim-to-real transfer technique, namely randomization of muscle dynamics, is adopted to bridge the gap between simulation and real-world deployment. Extensive experiments and ablation studies are conducted utilizing two string-type artificial muscle-driven robotic systems including a two degree-of-freedom robotic eye and a parallel robotic wrist, the results of which demonstrate the effectiveness of the proposed learning control strategy.
Index Terms:
Machine learning for robot control, artificial muscle, deep reinforcement learning, sim-to-real.
I Introduction
Artificial muscles, designed to mimic biological counterparts, generate mechanical forces and movements in response to external stimuli like electrical or chemical signals. Well known for their high power-to-weight ratios and inherent compliance [1], artificial muscles offer a unique combination of strength and flexibility conducive to versatile and adaptive robot designs. Over the past decade, robotic systems driven by artificial muscles have achieved significant advancements, demonstrating impressive performance across diverse domains, including biomimetic robots [2, 3, 4], robot manipulators and grippers [5, 6], soft robots [7, 8], and many others [9, 10, 11].
Despite their potential, achieving accurate control of string-type artificial muscle-driven robots remains a challenge due to their highly nonlinear and complex dynamics. While previous research efforts have primarily centered around structural design and open-loop motion performance evaluation [12, 13, 14], recent studies have begun to investigate closed-loop control strategies with a majority exploring proportional-integral-derivative (PID) controllers [15, 16, 17]. However, PID controllers often fall short in consistently delivering desired control performance, especially in complex robotic systems.
With the rapid development of machine learning techniques, learning-based controllers have emerged as an increasingly popular avenue in the muscle-driven robotics community. For instance, N. Nikdel et al. [6] developed a neural network-based model predictive controller for robotic manipulators actuated by shape memory alloys (SMAs). M. Giorelli et al. [18] established the inverse kinematics of cable-driven soft robotic arms using feedforward neural networks. T. Yang et al. [19] proposed a reinforcement learning control strategy for a cuttlefish-like soft robot driven by dielectric elastomer (DE) membranes. Among these approaches, deep reinforcement learning (DRL) stands out for its remarkable performance across diverse and complex tasks, offering a universal control solution. An increasing number of researchers have leveraged DRL algorithms to address control challenges in artificial muscle-driven robots, spanning applications ranging from dexterous grasping of robotic hands[20] to agile motion control of soft robotic manipulators [21, 22], and various others [23, 24, 25, 26].
In most of existing studies [20, 22, 24, 23, 26], the standard procedure of applying DRL control in the targeted robotic systems typically follows the schematic illustrated in Fig. 1. Initially, it is crucial to identify the relevant parameters of the artificial muscles and formulate the system dynamics model to establish a simulation environment. Subsequently, the DRL algorithms undergo training within simulation to develop a control policy. Finally, the acquired control policy is deployed on the robotic platform to accomplish various control objectives. Despite these advancements, challenges persist in the application of DRL control in the targeted robotic systems.
Figure 1: Schematic of standard procedure of applying DRL control in artificial muscle-driven robotic systems. Three main steps include 1) dynamic modeling with system identification to establish simulation environment; 2) training DRL agent within the environment to develop a control policy; and 3) deploying the trained policy onto real-world robotic systems for testing.
I-1 Muscle Sim-to-real Gap
For the robotic systems driven by artificial muscles, modeling errors primarily come from the following three aspects. First, the dynamic models of artificial muscles themselves are approximate and inaccurate [27, 28]. Second, inevitable errors occur in the system identification process of the muscles. Third, different muscle strings typically have different model parameters [29, 30], and conducting system identification for all muscle strings is time-consuming and labor-intensive. These factors jointly contribute to inaccurate modeling of artificial muscles, resulting in a significant sim-to-real gap for the entire robotic system. Consequently, control strategies that perform well in simulation may not seamlessly transfer to real-world applications [21, 22]. Despite being labor-intensive, some researchers [19, 25] have attempted to directly train DRL algorithms on real robots. However, given the fragile nature of string-type artificial muscles [1], this approach lacks universality and safety guarantees. Hence, addressing the muscle sim-to-real gap for DRL-based controllers is imperative.
I-2 Low Learning Efficiency
In reinforcement learning, an agent continually interacts with the environment to acquire a good policy through learning from the interaction data [31]. Due to the complexity and high-dimensional nature of deep neural networks, DRL algorithms usually require a large amount of training data and iterations for convergence towards the desired policy. While training in simulation remains feasible, many studies have noted that the low sample efficiency often leads to significantly prolonged training times [23, 26, 32]. Moreover, the complex dynamic models of muscle-driven robots exacerbate this issue. Thus, enhancing training data efficiency and accelerating convergence speed are of great importance.
We observe a scarcity of research addressing the aforementioned issues within the domain of artificial muscle-driven robots. Our prior research [24] aimed to tackle the challenge of low learning efficiency. However, its experimental findings were solely derived from simulations, without addressing the crucial muscle sim-to-real gap. This paper presents a DRL-based control strategy with improved data efficiency and transferability to real-world applications. It is worth noting that in recent years, there have been some efforts to bridge the sim-to-real gap of artificial muscles. Many researchers [27, 28, 33, 34, 35] have proposed more accurate yet significantly more complex dynamic models. In this paper, however, we seek to directly improve the performance of DRL-based controllers by employing a straightforward sim-to-real transfer technique, thereby bypassing the complicated process of mathematical modeling and parameter identification for artificial muscle strings. We evaluate our learning control frame work on two robotic systems powered by artificial muscle strings, including a biomimetic robotic eye driven by super-coiled polymers (SCPs) [36, 29], and a parallel robotic wrist actuated by twisted-coiled actuators (TCAs) [30, 37].
The main contributions of this paper are as follows. To the best of our knowledge, this paper proposes the first efficient learning control framework designed for string-type artificial muscle-driven robotic systems. Specifically, we introduce a novel augmentation method for state-based data, which transforms a single transition data into several. We leverage an easy-to-implement PID controller to generate trajectories for early-stage training of DRL agents. The training results in simulation show that our proposed methods improve the data efficiency of the DRL algorithm in targeted robotic systems. Additionally, we leverage a zero-shot sim-to-real transfer technique on artificial muscle dynamics. The experimental results demonstrate a great improvement in real-world control performance through the application of this method in muscle-driven robotic systems. We believe that the proposed learning control framework can be utilized to address a variety of control tasks across similar robotic systems.
The remainder of this paper is organized as follows. Section II presents basic information about leveraging reinforcement learning within the robotic systems of interest. Section III offers a concise overview of the robotic platforms utilized in this study. Section IV presents our proposed efficient learning control framework in detail. Subsequently, Section V presents the experimental results of the proposed method, encompassing training outcomes in simulation and control performances in real-world scenarios. Finally, concluding remarks and current limitations are provided in Section VI.
II Preliminaries
In this section, we model the control problem of targeted robotic systems as a reinforcement learning problem by Markov Decision Process (MDP), which is a mathematical framework that describes the interaction between the DRL agent and the environment [31]. The MDP in artificial muscle-driven robots consists of the following tuple, i.e.,
$(\mathcal{S},\mathcal{A},\mathcal{R},f,\gamma).$ (1)
  • $\mathcal{S}$ is the state space. At each time step $t$, the state vector $s_{t}\in\mathcal{S}$ generally includes the current state variables of the robotic system and the target positions and/or orientations. In this paper, the motion state variables of the robotic systems are collectively defined as vector $\bm{x}$, which contains the Euler angles, angular velocities, and other relevant motion states. The system output vector, containing the variables to be controlled, is denoted by $\bm{y}$. Correspondingly, the target system output is denoted by $\bm{y}^{*}$. In this way, the state vector $s_{t}$ is represented as
    $s_{t}=(\bm{x}_{t},\bm{y}^{*}).$ (2)
  • $\mathcal{A}$ is the action space, where each action $a_{t}\in\mathcal{A}$ is generated by the DRL agent based on the current state $s_{t}$ and its own policy. The action vector typically consists of the actuating voltage applied to each muscle or the voltagel difference between a pair of antagonistic muscles.
  • $\mathcal{R}$ is the reward function that maps the current state $s_{t}$ and action $a_{t}$ to an immediate reward $r_{t}$, i.e.,
    $r_{t}=\mathcal{R}(s_{t},a_{t}).$ (3)
    The design of reward function plays a vital role in the learning process of the DRL agents, which learn the optimal policy under the guidance thereby. In this paper, we develop a generalized quadratic-form reward function design for the robotic systems of interest:
    $r_{t}=-(e^{T}_{t}Q_{e}e_{t}+a^{T}_{t}R_{a}a_{t})+r_{\text{bonus}}$ (4)
    where $e_{t}=|\bm{y}^{*}-\bm{y}_{t}|$ is the system tracking error, and $Q_{e}$ and $R_{a}$ represent the output and control weight matrices, respectively. A bonus reward $r_{\text{bonus}}$ is given when the system tracking error achieves a sufficiently small threshold. We add this bonus reward because, for the robotic systems discussed in this paper, maintaining at the target state is generally more challenging than reaching it.
  • $f$ represents the state transition function, which describes the likelihood of a transition from state $s$ to $s^{\prime}$ when an action $a$ is taken.
  • $\gamma\in[0,1]$ is the discount factor, which serves to balance the influence of future rewards on the current state. In all experiments conducted in this paper, we set $\gamma$ to a constant value of $\gamma=0.99$.
III Experimental Platforms
This section gives a concise overview of the testing platforms utilized in this paper, including an SCP-driven robotic eye (see Fig. 2) and a TCA-driven parallel robotic wrist (see Fig. 3). We briefly describe the hardware design, state space, action space, and reward function for each system.
Figure 2: The 2-DOF robotic eye prototype driven by two pairs of SCP muscles.
III-A Robotic Eye
The high occurrence of unsuccessful corrective surgeries for ocular motor disorders underscores the need for a muscle-driven robotic eye platform to assist ophthalmologists in gaining deeper insights into the biomechanics of human eye movement. In our previous work [36], we proposed a prototype of a 2-DOF robotic eye, and derived the geometric and dynamic models specific to this prototype. As illustrated in Fig. 2, the hemispherical eyeball is driven by two pairs of SCP muscle actuators, accommodating a Raspberry Pi Mini Camera (RPiMC). The primary control objective is to precisely and rapidly adjust the eyeball’s orientation towards the desired target angles by regulating the SCP muscles, namely foveation control in the eye movement study.
The eyeball’s center of rotation, denoted as $\bm{O}\in\mathbb{R}^{3}$, remains fixed at the origin within the reference frame $O_{xyz}$. Here, the axes $x$, $y$, and $z$ correspond to the transverse, sagittal, and vertical axes of the eyeball, respectively. Rotations around these axes are expressed as pitch $\theta$, roll $\phi$, and yaw $\psi$. The robotic eye system’s state variables are represented as
$\bm{x}=[\theta,\dot{\theta},\psi,\dot{\psi},\Delta T_{1},\Delta T_{2},\Delta T_{3},\Delta T_{4}]$ (5)
where $\theta,\dot{\theta},\psi,\dot{\psi},$ represent the pitch, yaw angles, and their respective angular velocities for control. Additionally, $\Delta T_{i}$ represents the temperature difference between the ambient temperature $T_{0}$ and the current temperature $T_{i}$ of muscle $m_{i}$.
Figure 3: The parallel robotic wrist prototype actuated by three TCAs.
Focusing primarily on the robotic eye’s orientation, the system’s output vector is denoted as $\bm{y}=[\psi,\dot{\psi},\theta,\dot{\theta}]$. Due to the difficulty and poor accuracy of measuring the temperature of each muscle in real time, the state vector of the DRL agent is defined by
$s_{t}=[\theta_{t},\dot{\theta}_{t},\psi_{t},\dot{\psi}_{t},\theta^{*},\psi^{*}].$ (6)
Here, $\theta^{*}$ and $\psi^{*}$ represent the target angles. The Euler angles and corresponding angular velocities are measured through visual feedback from the embedded RPiMC. The actuating voltage applied to muscle $m_{i}$ is denoted as $V_{i}$. Given the antagonistic configuration of the muscle pairs, the control input is defined by a constrained action vector, i.e.,
$a_{t}=[$ $-V_{1}+V_{2},-V_{3}+V_{4}]$ (7)
s.t. $V_{1}\cdot V_{2}=V_{3}\cdot V_{4}=0$
$0\leq V_{i}\leq 10$
The constraint requires that only one muscle is activated at each time step in the pairs $(m_{1},m_{2})$ and $(m_{3},m_{4})$. The reward function follows the structure of (4), where the output and control weight matrices are set to $Q_{e}=\text{diag}\{0.05,0.25,0.05,0.25\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively. The bonus reward is granted when the tracking error is sufficiently small, given by
$r_{\text{bonus}}=2\times\left(\mathds{1}\{|\psi-\psi^{*}|<0.3\}+\mathds{1}\{|\theta-\theta^{*}|<0.3\}\right).$ (8)
Here, $\mathds{1}\{\cdot\}$ denotes the binary indicator function that yields a value of $1$ if the given condition holds true, and 00 otherwise.
Figure 4: Overview of the proposed efficient learning control framework, including a two-stage training process in simulation and real-world application. Stage 1 utilizes bootstrapping with PID control, aiming at enhancing data efficiency. Stage 2 integrates the baseline SAC learning algorithm with state-based data augmentation and randomization of muscle dynamics for expediting training convergence and bridging sim-to-real gap, respectively.
III-B Robotic Wrist
Research into artificial muscle-driven robotic wrist joints holds considerable importance in several domains, including the development of dexterous robotic systems with human-like capabilities, advancements in rehabilitation and prosthetics technology, and the enhancement of robotic manipulation versatility across various fields. We designed a parallel robotic wrist prototype driven by three equally spaced compound TCAs [30] as illustrated in Fig. 3. The base plate remains stationary, while the upper plate, also known as the end effector, moves via three muscle strings. Linkages between the two plates are constructed to support two degree-of-freedom motion of the upper plate [38]. Given this configuration, we developed a simplified dynamics model under the assumption of a fixed rotational center for the upper plate. Similar to the foveation control of the robotic eye, the control objective entails precise regulation of the moving upper plate’s orientation. This necessitates adjusting the applied actuating voltages across the three muscle strings to drive the upper plate from its initial equilibrium state to predefined target angles.
The reference frame $O_{xyz}$ is fixed on the moving upper plate, with the $z$-axis perpendicular to the upper plate and pointing outward. $\theta,\phi,$ and $\psi$ represent the pitch, roll, and yaw angles, respectively. Similar to the robotic eye, the system state variables are defined as
$\bm{x}=[\theta,\dot{\theta},\phi,\dot{\phi},\Delta T_{1},\Delta T_{2},\Delta T_{3}]$ (9)
where $\dot{\theta}$ and $\dot{\phi}$ represent the corresponding angular velocities, and $\Delta T_{i}$ is the temperature difference of muscle $m_{i}$ with respect to the ambient environment, with $i\in\{1,2,3\}$. The system output vector is represented as $\bm{y}=[\theta,\dot{\theta},\phi,\dot{\phi}]$. Likewise, the muscle’s temperature is not used as an input for the DRL agent, and thus, the state vector is defined as
$s_{t}=[\theta_{t},\dot{\theta}_{t},\phi_{t},\dot{\phi}_{t},\theta^{*},\phi^{*}].$ (10)
Here, $\theta^{*}$ and $\phi^{*}$ denote the target angles, with the respective Euler angles and angular velocities measured using the inertial measurement unit (IMU) positioned at the center of the upper plate. Additionally, the input action vector $a_{t}$ of the system is defined as
$a_{t}=[V_{1},V_{2},V_{3}]$ (11)
where $V_{i}$ represents the actuating voltage applied to muscle $m_{i}$, constrained by $0\leq V_{i}\leq 10$. The output and control weight matrices of the reward funtion are set to $Q_{e}=\text{diag}\{0.05,0.2,0.05,0.2\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively. When the tracking error is sufficiently small, a bonus reward is granted, i.e.,
$r_{\text{bonus}}=2\times\left(\mathds{1}\{|\theta-\theta^{*}|<0.5\}+\mathds{1}\{|\phi-\phi^{*}|<0.5\}\right).$ (12)
IV Learning Control Framework
This section provides a comprehensive overview of the proposed learning control framework, as illustrated in Fig. 4. The soft actor-critic (SAC) serves as the baseline DRL algorithm. To enhance data efficiency, we employ a two-stage training strategy. Initially, a standard PID controller is initialized to generate imperfect trajectories for the DRL agent. Subsequently, the agent interacts directly with the simulation environment to update its policy. All transition data observed from the environment undergoes an state-based augmentation strategy before being stored into the replay buffer. To address the sim-to-real gap, we introduce a randomization method on the artificial muscle dynamics at the beginning of each training episode. Eventually, the trained actors (control policies) are able to be directly applied in real-world robotic systems to achieve desired control performance.
IV-A Soft Actor-Critic (SAC)
SAC, a state-of-the-art deep reinforcement learning algorithm, finds extensive application across various tasks [39]. Traditional DRL algorithms primarily maximize cumulative rewards, i.e.,
$J(\pi)=\sum_{t=0}^{T}\mathbb{E}_{(s_{t},a_{t})\sim\rho_{\pi}}[\mathcal{R}(s_{t},a_{t})]$ (13)
where $\pi$ represents the policy. However, SAC adopts an entropy-maximizing approach, with its learning objective formulated as
$J(\pi)=\sum_{t=0}^{T}\mathbb{E}_{(s_{t},a_{t})\sim\rho_{\pi}}[\mathcal{R}(s_{t},a_{t})+\alpha\mathcal{H}(\pi(\cdot|s_{t}))].$ (14)
where $\mathcal{H}(\pi(\cdot|s_{t}))$ represents an entropy of the policy $\pi$, and $\alpha$ serves as a temperature parameter. This approach enhances the robustness of DRL and accelerates its training process [31].
Figure 5: Schematic diagram of the network architecture. The actor (left) and the critic (right) networks share the same structure but have different parameters. Each is composed of a fully-connected layer with ReLU activation as the input layer, a recurrent GRU [40] unit, and a fully-connected output layer.
In addition, SAC is an off-policy algorithm that leverages the actor-critic architecture. The neural network structures utilized in this paper are illustrated in Fig. 5. To process the sequential information crucial for implementing dynamics randomization and handling partial observability, we employ a Gated Recurrent Unit (GRU) [40]. We select GRU over over commonly-used Long Short-Term Memory (LSTM) due to its efficiency and ease of training [40]. Noise is added to the observation input to the actor. The hyperparameters of the SAC algorithm are detailed in Table I and remain consistent across all ablation studies conducted in this paper.
TABLE I: SAC Hyper-parameters
Parameter Value
Learning rate of actor $1\times 10^{-4}$
Learning rate of critic $1\times 10^{-4}$
Replay buffer size $1\times 10^{5}$
Mini-batch size $20$
Entropy target $-\text{dim}(\mathcal{A})$ (e.g., $-3$ for robotic wrist)
Soft update factor $\tau$ $0.005$
Discount factor $\gamma$ $0.99$
Hidden layer size of GRU $256$
Gradient steps $k$ $40$
Target networks update interval $1$
IV-B Randomization of Muscle Dynamics
The core idea behind dynamics randomization is to introduce randomness into specific dynamic parameters within the simulator during training. This approach allows the DRL agent to learn a policy adaptable to various dynamics conditions, including those encountered in real-world experiments. Widely adopted in diverse robotic applications [41, 42, 43], this method serves as inspiration for our work. Specifically, we randomize several critical parameters within the simulated dynamics environment. Among the robotic systems discussed in this paper, the dynamic parameter set of the muscles, denoted by $\mu$, holds pivotal importance. Therefore, we establish a distribution $\rho_{\mu}$ to encompass the variability of muscle dynamic parameters for dynamics randomization purposes. Additionally, we incorporate Gaussian noise into the observation inputs of the policy to emulate the noise commonly present in real-world sensor data.
Regarding the SCP muscles driving the robotic eye and the TCA muscles driving the parallel robotic wrist, their thermomechanical models are approximately as [29]
$F=k(x-x_{0})+b\dot{x}+c(T(t)-T_{\text{amb}}).$ (15)
Here, $x$ and $x_{0}$ represent the loaded length and unloaded length of the actuator, respectively. $k$, $b$, and $c$ denote the mean stiffness, damping, and temperature effect of the actuator, respectively. $T(t)$ denotes the current temperature, while $T_{\text{amb}}$ stands for the ambient temperature of the environment. The thermoelectric model of the actuator is expressed as
$C_{\text{th}}\frac{dT(t)}{dt}=P(t)-\lambda(T(t)-T_{\text{amb}}).$ (16)
Here, $C_{\text{th}}$ represents the thermal mass of the actuator, and $\lambda$ represents the absolute thermal conductivity of the actuator. $P(t)$ stands for the heat power applied to the actuator, calculated as $P(t)=V^{2}/R$, where $V$ is the voltage applied to the actuator and $R$ is the electrical resistance. In summary, the dynamic parameter set for randomization is
$\mu:=\{k,b,c,C_{\text{th}},\lambda,R\}$ (17)
TABLE II: Dynamics parameters of randomization and their respective range in the robotic eye
Parameter Value Scailing factor Additive term
$k$ $(N/cm)$ $0.25$ uniform$(0.8,1.2)$
$b$ $(N\cdot s/cm)$ $0.01$ uniform$(0.9,1.1)$
$c$ $(N/^{\circ}C)$ $0.0055$ uniform$(0.85,1.15)$
$C_{\text{th}}$ $(Ws/^{\circ}C)$ $0.28$ uniform$(0.8,1.2)$
$\lambda$ $(W/^{\circ}C)$ $0.094$ uniform$(0.85,1.15)$
$R$ $(\Omega)$ $20.0$ uniform$(0.9,1.1)$
$\theta,\psi$ noise $(^{\circ})$ normal$(0,0.1)$
$\dot{\theta},\dot{\psi}$ noise $(^{\circ}/s)$ normal$(0,0.05)$
Table II and III outline the distribution $\rho_{\mu}$ of SCP and TCA muscles, respectively, along with the observation noise added for the robotic eye and wrist systems. $\rho_{\mu}$ is centered on the mean values derived from system identification to align the data distribution in simulation more accurately with that in reality. The range of randomization for each parameter is determined based on the distribution observed across different real muscle strings, as outlined in previous work [29]. During the training process, the physical parameters of the actuators are randomized within the scaling ranges (i.e., $\mu\sim\rho_{\mu}$) at the beginning of each training episode and remain constant throughout. Whereas, the observation noise varies at each time step and across different Euler angles. While this paper employs a simplified linear dynamic model for artificial muscles, this method is expected to be readily transferred to more complex muscle dynamics models.
IV-C Bootstrap via PID
On-line DRL agents mainly rely on random exploration [31]. However, due to the slow response of artificial muscles, random exploration policies may struggle to explore the state space comprehensively and rapidly, resulting in low sample efficiency. To address this issue, we draw inspiration from previous work [44, 45] and boost the training process by providing demonstrations into initial replay buffer. While PID controllers demand extensive tuning for optimal control performance, their implementation is relatively straightforward, given that moderate performance suffices. In the case of muscle-driven robotic systems, PID policies typically suffice to approach the target state. Hence, they provide trajectories that explore the defined state space more comprehensively, serving as an advantageous starting point for the DRL learning process. Subsequently, the agent will interact with the environment and continue to refine its policy, thus eventually outperforming PID and achieving desired control performances.
Before the DRL training process begins, we initialize a PID controller that offers a barely acceptable (and typically not excellent) control performance. Within the first $M$ episodes of training, the PID controller interacts with the environment and generates trajectories, which are then stored in the replay buffer along with the the rewards. Differing from traditional behavior cloning methods, we don’t utilize a supervised loss during demonstration learning. Consequently, the agent does not strictly follow the PID control policy but leverages the trajectories generated by PID solely to bootstrap the initial training.
IV-D Augmentation of State Vector
Data augmentation refers to a class of techniques aimed at enhancing the size and quality of training datasets, thereby facilitating the development of more robust deep learning models [46]. The fundamental concept of data augmentation involves transforming limited data into diverse, functionally equivalent datasets. Given the labor-intensive and computationally expensive nature of collecting training data in robotic systems, we integrate data augmentation into the agent’s learning process. While prior research [47, 48, 49] has explored data augmentation in reinforcement learning, existing studies have primarily focused on pixel-based RL scenarios, where agents rely on image-based observations. In this paper, we present a straightforward and adaptable approach to state-based data augmentation, better suited and more effective for our targeted robotic systems.
TABLE III: Dynamics parameters of randomization and their respective range in the robotic wrist
Parameter Value Scailing factor Additive term
$k$ $(N/mm)$ $0.21$ uniform$(0.8,1.2)$
$b$ $(N\cdot s/cm)$ $0.63$ uniform$(0.9,1.1)$
$c$ $(N/^{\circ}C)$ $0.0707$ uniform$(0.85,1.15)$
$C_{\text{th}}$ $(Ws/^{\circ}C)$ $3.06$ uniform$(0.8,1.2)$
$\lambda$ $(W/^{\circ}C)$ $0.1189$ uniform$(0.85,1.15)$
$R$ $(\Omega)$ $10.0$ uniform$(0.9,1.1)$
$\theta,\phi$ noise $(^{\circ})$ normal$(0,0.1)$
$\dot{\theta},\dot{\phi}$ noise $(^{\circ}/s)$ normal$(0,0.05)$
As commonly acknowledged, rewards serve as pivotal guides in the learning process of DRL agents. Therefore, our objective is to generate more data transitions with different rewards. By combining (2) and (3), the reward $r_{t}$ is then calculated as
$r_{t}=\mathcal{R}(\bm{x}_{t},\bm{y}^{*},a_{t}).$ (18)
Here, $\bm{x}_{t}$ and $a_{t}$ are directly related to system dynamics, whereas the target output vector $\bm{y}^{*}$ is manually set and changeable. Consequently, the proposed augmentation of state vector proceeds as follows. At each time step $t$, we fix the motion state vectors $\bm{x}_{t},\bm{x}_{t+1}$ and the action $a_{t}$, and vary the target vector $\bm{y}^{*}$. In this paper, we introduce a random additive term into $\bm{y}^{*}$ as follows
$\bm{y}^{*}_{\text{new}}=\bm{y}^{*}\pm\delta\bm{Z}.$ (19)
Here, $\bm{y}^{*}_{\text{new}}$ denotes the new target vector. $\delta$ acts as a scalar factor to control the scope of augmentation, which largely depends on the defined range. $\bm{Z}$ is a random vector sharing identical dimensions with $\bm{y}^{*}$, wherein the elements of $\bm{Z}$ range from 00 to $1$. Thus, we derive a pair of new states $s^{\prime}_{t}$ and $s^{\prime}_{t+1}$. After that, a new corresponding reward $r^{\prime}_{t}$ is calculated using (18). Following the above steps, the original transition tuple is transformed into a new one, i.e.,
$(s_{t},a_{t},r_{t},s_{t+1})\xrightarrow{AUG}(s^{\prime}_{t},a_{t},r^{\prime}_{t},s^{\prime}_{t+1}).$ (20)
Algorithm 1 Augmentation of State-based Data
1:  Generate rollout $\tau=(s_{0},a_{0},r_{0},...,s_{T})$ from environment
2:  for $i=1$ to $n$ do
3:     for each $s_{t},a_{t},r_{t}$ in $\tau$ do
4:        Create a new state vector $s_{t}\rightarrow s^{\prime}_{t}$
5:        Calculate a new reward $r^{\prime}_{t}=\mathcal{R}(s^{\prime}_{t},a_{t})$
6:     end for
7:     Get a new rollout $\tau^{\prime}_{i}=(s^{\prime}_{0},a_{0},r^{\prime}_{0},...,s_{T})$
8:  end for
TABLE IV: Simulation Parameters of Robotic Systems
Description Symbol Value
Robotic Eye Total time duration of one episode $t_{\text{total}}$ $15$s
Action time step $t_{a}$ $0.5$s
Total number of training episodes $N$ $2000$
Random or boost episodes $M$ $250$
Absolute value of $\theta$ bound $\theta_{\text{max}}$ $15^{\circ}$
Absolute value of $\psi$ bound $\psi_{\text{max}}$ $10^{\circ}$
Robotic Wrist Total time duration of one episode $t_{\text{total}}$ $30$s
Action time step $t_{a}$ $0.5$s
Total number of training episodes $N$ $3500$
Random or boost episodes $M$ $500$
Absolute value of $\theta$ bound $\theta_{\text{max}}$ $15^{\circ}$
Absolute value of $\phi$ bound $\phi_{\text{max}}$ $15^{\circ}$
Algorithm 2 SAC-BAR Learning Control
1:  Initialize critic networks $Q_{\theta_{1}},Q_{\theta_{2}}$, and actor network $\pi_{\phi}$ with random parameters $\theta_{1},\theta_{2},\phi$
2:  Initialize target network parameters $\bar{\theta}_{1}\leftarrow\theta_{1},\bar{\theta}_{2}\leftarrow\theta_{2}$
3:  Initialize an empty replay buffer $\mathcal{B}$
4:  Initialize a PID controller $\pi_{\text{PID}}$
5:  Initialize dynamics parameters distribution $\rho_{\mu}$
6:  for $\text{step}=1$ to $M$ do
7:     Sample dynamics $\mu\sim\rho_{\mu}$
8:     Generate rollout $\tau_{0}=(s_{0},a_{0},r_{0},...,s_{T})$ with dynamics $\mu$ and $a_{t}\sim\pi_{\text{PID}}(s_{t})$
9:     for $i=1$ to $n$ do
10:        Create new rollouts $\tau_{0}\xrightarrow{AUG}\tau_{i}$
11:     end for
12:     Store $\tau_{0},\tau_{1},...,\tau_{n}$ into $\mathcal{B}$
13:  end for
14:
15:  for $\text{step}=M+1$ to $N$ do
16:     Sample dynamics $\mu\sim\rho_{\mu}$
17:     Generate rollout $\tau_{0}=(s_{0},a_{0},r_{0},...,s_{T})$ with dynamics $\mu$ and $a_{t}\sim\pi_{\phi}(s_{t})$
18:     for $i=1$ to $n$ do
19:        Create new rollouts $\tau_{0}\xrightarrow{AUG}\tau_{i}$
20:     end for
21:     Store $\tau_{0},\tau_{1},...,\tau_{n}$ into $\mathcal{B}$
22:     for $i=1$ to $k$ do
23:        Sample a mini-batch of rollouts from $\mathcal{B}$
24:        Calculate the gradients $\nabla J_{Q}(\theta_{i}),\nabla J_{\pi}(\phi),\nabla J(\alpha)$
25:        Update network weights:
26:        $\theta_{i}\leftarrow\theta_{i}-\lambda_{Q}\nabla J_{Q}(\theta_{i})$
27:        $\phi\leftarrow\phi-\lambda_{\pi}\nabla J_{\pi}(\phi)$
28:        $\bar{\theta}_{i}\leftarrow\tau\theta_{i}+(1-\tau)\theta_{i}$
29:        Adjust temperature parameter:
30:        $\alpha\leftarrow\lambda_{\alpha}\nabla J(\alpha)$
31:     end for
32:  end for
The procedure of using augmentation for a state-based trajectory $\tau$ is provided in Algorithm 1. Each time the agent interacts with the environment and gets a new trajectory, this method creates $n$ different trajectories at a very low computational cost. We set $n=10$ in both robotic eye and wrist systems during training. In this paper, we denote the bootstrap method by the letter ’B’, augmentation method by the letter ’A’, and randomization of muscle dynamics by the letter ’R’. Algorithm 2 provides a detailed overview of the entire training procedure for the SAC algorithm with all three proposed improvements (i.e., SAC-BAR).
V Experimental Results
This section presents the experimental results of the proposed efficient learning control framework within two testing robotic platforms. We first present the training results of the algorithm in simulation environments, validating the effectiveness of the proposed bootstrap and augmentation methods in enhancing training data efficiency. Subsequently, we demonstrate the results of an ablation study conducted on the robotic wrist platform, verifying the effectiveness of dynamics randomization in improving the real-world control performance. Finally, experimental results on robotic eye are provided to demonstrate the generalizability of the proposed control framework across different robotic systems of interest.
(a) Robotic Eye
(b) Robotic Wrist
Figure 6: Reward trajectories with respect to the training episode number in ablation study. Ours refers to the proposed SAC-BAR algorithm, where B, A, and R indicate bootstrap, augmentation, and randomization techniques, respectively. Performance is evaluated in simulation with random dynamics every $10$ training episodes. The shaded area indicates the standard deviation of the mean reward based on running each experiment with three different random seeds.
Figure 7: Experimental trajectories of vanilla SAC algorithm on the robotic wrist under target angles ($\theta^{*}=-5^{\circ},\phi^{*}=7.5^{\circ}$) and ($\theta^{*}=2.5^{\circ},\phi^{*}=-7.5^{\circ}$), respectively. (a), (c) The system output angles $(\theta,\phi)$. (b), (d) The corresponding actions $a_{t}$ generated by the trained actor.
Figure 8: Experimental trajectories of our proposed SAC-BAR algorithm on the robotic wrist under target angles ($\theta^{*}=-5^{\circ},\phi^{*}=7.5^{\circ}$) and ($\theta^{*}=2.5^{\circ},\phi^{*}=-7.5^{\circ}$), respectively. (a), (c) The system output angles $(\theta,\phi)$. (b), (d) The corresponding actions $a_{t}$ generated by the trained actor.
V-A Training Results
We established simulation environments for two robotic systems to train DRL agents and conduct preliminary tests. Table IV lists the simulation parameters for each robotic system. The overall duration of a single episode $t_{\text{total}}$ and the action time step $t_{a}$ depend on the rise time and response speed constraints inherent to the hardware of the systems. The values of $N$ and $M$ are empirical parameters contingent on the complexity of the systems. For simpler dynamics, smaller values for $N$ and $M$ facilitate the learning process for the DRL agent in acquiring an effective control policy. For algorithms not using the bootstrap method, random exploration is employed in the first $N$ episodes. Target angles are confined within the motion range of the robotic device. With the majority of training time spent on dynamics calculation, the time step for dynamics simulation is set at $0.01$s to balance simulation accuracy and computational efficiency. Training and testing in simulation are conducted on a computer with an Intel Xeon Platinum 8383C CPU and an NVIDIA GeForce RTX 4090 GPU. The algorithms are implemented in PyTorch and trained using the Adam optimizer.
In each training episode, a target state is randomly selected from a uniformly distributed range. For instance, in the robotic eye scenario, the target angles adhere to $\theta^{*}\sim\mathcal{U}[-15^{\circ},15^{\circ}]$ and $\psi^{*}\sim\mathcal{U}[-10^{\circ},10^{\circ}]$. The robotic agent endeavors to reach the target from the initial state throughout the episode. For algorithms incorporating dynamics randomization, a new set of muscle dynamics parameters $\mu$ is sampled at the beginning of each episode by drawing values for each parameter from their respective range. Additionally, simulated sensor noise is introduced to the observed state at every action step. In the absence of dynamics randomization, muscle parameters are kept constant at their mean values, and no simulated noise is added during the training process.
To evaluate the effectiveness of the proposed bootstrap and augmentation techniques, we conducted an ablation study on two robotic systems. Training results of our learning control design, with and without bootstrap and augmentation, are presented in Fig. 6. The policies were trained and evaluated in simulation with random dynamics. As illustrated in the figures, our proposed algorithm exhibits the highest level of data efficiency in both robotic systems. Compared to the SAC algorithm trained solely with dynamics randomization (referred to as ”No BA”), the integration of bootstrap and augmentation enhances data efficiency and expedites the training process. In the robotic eye system, the baseline algorithm achieves an average reward of $-0.10$ after $2000$ training episodes, whereas our approach achieves comparable performance in only approximately $750$ episodes, resulting in a data efficiency increase to $2.67$ times the original level. Similarly, in the robotic wrist system, the baseline algorithm achieves an average reward of $-1.61$ after $3500$ episodes, while ours achieves comparable performance in only about $1200$ episodes, leading to a data efficiency increase to $2.92$ times. Notably, in the case of the more complex robotic wrist system, the demonstrations provided by the PID controller yield a more significant acceleration effect during the early stages of training. Meanwhile, the augmentation method demonstrates a favorable impact on enhancing the training speed in both systems.
(a) vanilla-SAC
(b) SAC-R
(c) SAC-BAR
(d) Box-and-Whisker Plot
Figure 9: Experimental results of the steady-state error tests in the parallel robotic wrist under selected control algorithms. (a), (b), (c) Heat maps illustrating the experimental results for the vanilla SAC algorithm, SAC with the randomization of muscles only (SAC-R), and our complete efficient learning control algorithm (SAC-BAR), respectively. The combination of target angles $(\theta^{*},\phi^{*})$ is systematically spaced at $2.5^{\circ}$ intervals within the range $(|\theta^{*}|,|\phi^{*}|)\leq 10^{\circ}$. The color bars in the visualization indicate the Euclidean distance norm of steady-state errors. (d) The box charts illustrating the medians, maxima, minima, quartiles, and outliers of the steady-state errors for the selected comparative control algorithms.
V-B Ablation Study in Robotic Wrist Experiments
To evaluate the impact of each component on real-world control performances, three different control policies are tested on the robotic wrist platform, including the vanilla SAC algorithm, SAC with only randomization of muscles (SAC-R), and our proposed efficient learning control framework (SAC-BAR). All three algorithms undergo training for an equivalent number of episodes in simulation, employing identical hyperparameters as specified in Table IV. During each experimental episode, the parallel robotic wrist starts from the initial state with $\theta=\phi=0^{\circ}$ and attempts to reach and maintain the target angle set $(\theta^{*},\phi^{*})$. The action vector $a$ is computed by the trained actors at a time step of $t_{a}=0.5$s, and each episode spans a duration of $t_{\text{total}}=20$s. Observations are measured by an IMU positioned at the center of the upper plate.
TABLE V: Averaged steady-state error over a total of 81 targets in ablation study with the robotic wrist
Agent Vanilla-SAC SAC-R SAC-BAR
Mean error (deg) $0.911\pm 0.457$ $0.657\pm 0.386$ $0.580\pm 0.368$
The experimental trajectories for two pairs of desired target angles, $(\theta^{*},\phi^{*})=(-5^{\circ},7.5^{\circ})$ and $(\theta^{*},\phi^{*})=(2.5^{\circ},-7.5^{\circ})$, of the vanilla SAC algorithm and our proposed SAC-BAR are illustrated in Fig. 7 and Fig. 8, respectively. In Fig. 7LABEL:sub@subfig:w_tra1_a and LABEL:sub@subfig:w_tra1_c, it is evident that without randomization of muscle dynamics, the vanilla-SAC agent struggles to achieve the desired orientation. While the agent may eventually reach the desired orientation at certain target angles, maintaining that target angle becomes a challenge, resulting in a considerable steady-state error. Conversely, as observed in Fig. 8LABEL:sub@subfig:w_tra2_a and LABEL:sub@subfig:w_tra2_c, the proposed SAC-BAR agent achieves the target state in less than $10$s and consistently remains within its neighborhood thereafter. This success is attributed to the agent learning to maximize the action value during the rising phase and promptly reduce it upon approaching the target angles, as illustrated by the input action trajectories in Fig. 8LABEL:sub@subfig:w_tra2_b and LABEL:sub@subfig:w_tra2_d. Additionally, when overshoot occurs, the agent still learns to return to and maintain the target orientation. Conversely, in Fig. 7LABEL:sub@subfig:w_tra1_b and LABEL:sub@subfig:w_tra1_d, there exist serious oscillations in the action trajectories of vanilla SAC agent during the initial $10$s, leading to a prolonged rise time and sometimes even failure to reach the target.
Figure 10: Experimental trajectories of the foveation test of the 2-DOF robotic eye. Target angles are set at $(\theta^{*}=-5^{\circ},\psi^{*}=7.5^{\circ})$ and $(\theta^{*}=10^{\circ},\psi^{*}=-5^{\circ})$, respectively. (a), (c) The system output angles $(\theta,\psi)$. (b), (d) The corresponding actions $a_{t}$ generated by the trained actor.
In order to comprehensively evaluate the control performance of the algorithms across different target angles, we design a steady-state error field test for the robotic wrist. We select a series of target angles $(\theta^{*},\phi^{*})$ spaced at $2.5^{\circ}$ intervals within the defined range $\theta^{*},\phi^{*}\in[-10^{\circ},10^{\circ}]$. Each agent is set to run for $25$s, sequentially targeting each angle pair. The steady-state error is measured and averaged over the last $5$s window, determined by computing the Euclidean distance norm between the current and target angles using $e_{\text{ss}}=\sqrt{(\theta-\theta^{*})^{2}+(\phi-\phi^{*})^{2}}$.
The steady-state errors of the field test of vanilla-SAC, SAC-R, and SAC-BAR are presented in Fig. 9LABEL:sub@subfig:a, LABEL:sub@subfig:b, and LABEL:sub@subfig:c, respectively. The $x$ and $y$ axes therein represent $\theta^{*}$ and $\phi^{*}$, while the color bars represent the averaged steady-state errors. Following that, box charts in Fig. 9LABEL:sub@subfig:d illustrate the medians, maxima, minima, quartiles, and outliers of the steady-state errors for different control policies. The figures demonstrate that vanilla-SAC agent performs poorly across nearly the entire defined range, whereas both SAC-R and SAC-BAR outperform significantly. The steady-state errors averaged over the defined range for three algorithms are presented in Table V. Compared to the baseline vanilla-SAC, SAC-R and SAC-BAR achieved reductions in averaged steady-state errors of $27.9\%$ and $36.3\%$, respectively. Specifically, our proposed SAC-BAR reaches the desired orientation at every target angle pair with steady-state errors consistently bounded below $1.5^{\circ}$ except for three outliers. The averaged steady-state error of SAC-BAR is $0.58^{\circ}$, demonstrating its superior control performance in accuracy and robustness across different target angles.
Figure 11: Experimental results in the steady-state error tests of the robotic eye at eyery combination of $(\theta^{*},\psi^{*})$ target angles spaced at $2.5^{\circ}$ intervals for $(|\theta^{*}|,|\psi^{*}|)\leq 10^{\circ}$ where the color bar represents the Euclidean distance norm of steady-state error.
The above experimental results indicate that randomization of muscle dynamics significantly enhances the real-world control performance of the learning-based controllers in the robotic system of interest, while bootstrap and augmentation techniques contribute less to the improvement. We believe that the enhancement observed in SAC-BAR compared to SAC-R primarily stems from the improvement in training efficiency. With the incorporation of bootstrap and augmentation methods, SAC-BAR achieves more comprehensive training and higher averaged rewards in simulation after $3500$ training episodes. This improvement is potentially more significant in scenarios where training data is limited or exploration proves more challenging.
V-C Learning Control of Robotic Eye
To demonstrate the applicability of the proposed learning control framework across diverse robotic systems driven by artificial muscle strings, this subsection presents experimental results on an additional platform, the 2-DOF robotic eye. We focus on the foveation control in the experiment. The robotic eye system, initialized at $\theta=\psi=0^{\circ}$ in ambient room temperature, is configured to foveate at various target angles $(\theta^{*},\psi^{*})$. Considering the system’s rise time, each testing episode lasts for $t_{\text{total}}=15$s, with an action time step of $t_{a}=0.5$s. The averaged Euclidean distance between the current and target angles within the last $5s$ window serves as the steady-state error metric, i.e., $e_{\text{ss}}=\sqrt{(\theta-\theta^{*})^{2}+(\psi-\psi^{*})^{2}}$.
Experimental trajectories with targets set at $(\theta^{*},\psi^{*})=(-5^{\circ},7.5^{\circ})$ and $(10^{\circ},-5^{\circ})$ are illustrated in Fig 10. These trajectories demonstrate that the measured angles rise to the target in approximately $5$s, with negligible steady-state error. A visual field test is conducted with a series of target angles spaced at $2.5^{\circ}$ intervals within the defined range $|\theta^{*}|,|\psi^{*}|\leq 10^{\circ}$ to evaluate the overall control performance of the algorithm. The robotic eye is tested to foveate at each target in sequence. The results, illustrated in Fig. 11, reveal excellent control performance, with steady-state errors consistently below $1^{\circ}$ for most target angles.
In prior research [23], we adopted a reinforcement learning-based controller using deep deterministic policy gradient (DDPG) algorithm [50], trained in simulation for $50000$ episodes. Additionally, a reduced-order state observer estimated the real-time temperature of SCP muscles, resulting in satisfactory control performance. In this paper, our proposed algorithm demonstrates comparable control performance with just $2000$ training episodes, representing only $1/25$ of the original requirement. Remarkably, it accomplishes this without the need for real-time muscle temperature information.
VI Conclusions
This paper presented a learning-based control framework designed for the string-type artificial muscle-driven robotic systems, leveraging the SAC algorithm with a GRU hidden layer as the baseline DRL algorithm. To enhance training data efficiency and expedite convergence speed, we proposed two methods including bootstrap by PID and state-based data augmentation. Additionally, we incorporated randomization of artificial muscle dynamic parameters during training to address the sim-to-real gap for enhancing real-world control performance. Extensive experiments and ablation studies on two systems—a 2-DOF robotic eye driven by SCP actuators and a parallel robotic wrist driven by TCA muscles—validated the effectiveness and generalizability of the proposed control framework.
In simulation, the designed bootstrap and augmentation methods significantly reduced the training episodes required for DRL-based controllers. Real-world application experiments demonstrated that randomization of muscle dynamics notably enhanced the control performance of learning-based controllers. Furthermore, our framework exhibited satisfactory performance on both robotic platforms, indicating its great potential of applications in similar systems driven by artificial muscle strings.
Despite the notable performance of the proposed algorithm, several limitations exist. Noticeable jitters in action trajectories may prolong system rise time. Additionally, the configuration of the reward function still has a considerable impact on the experimental outcomes. In future work, we plan to analyze these issues comprehensively using extensive experimental data and refine our algorithms to effectively address these challenges.

### v3

@@
 Index Terms:
-Machine learning for robot control, artificial muscle, deep reinforcement learning, sim-to-real.
+Reinforcement learning for robot control, artificial muscle, sim-to-real, sample efficiency.
 I Introduction
@@
 Well known for their high power-to-weight ratios and inherent compliance[1], artificial muscles offer a unique combination of strength and flexibility conducive to versatile and adaptive robot designs.
-Over the past decade, robotic systems driven by artificial muscles have achieved significant advancements, demonstrating impressive performance across diverse domains, including biomimetic robots [2, 3, 4], robot manipulators and grippers [5, 6], soft robots [7, 8], and many others [9, 10, 11].
-Despite their potential, achieving accurate control of string-type artificial muscle-driven robots remains a challenge due to their highly nonlinear and complex dynamics.
-While previous research efforts have primarily centered around structural design and open-loop motion performance evaluation [12, 13, 14], recent studies have begun to investigate closed-loop control strategies with a majority exploring proportional-integral-derivative (PID) controllers [15, 16, 17].
-However, PID controllers often fall short in consistently delivering desired control performance, especially in complex robotic systems.
-With the rapid development of machine learning techniques, learning-based controllers have emerged as an increasingly popular avenue in the muscle-driven robotics community.
-For instance, N.
-Nikdel et al.
-[6] developed a neural network-based model predictive controller for robotic manipulators actuated by shape memory alloys (SMAs).
-M.
-Giorelli et al.
-[18] established the inverse kinematics of cable-driven soft robotic arms using feedforward neural networks.
-T.
-Yang et al.
-[19] proposed a reinforcement learning control strategy for a cuttlefish-like soft robot driven by dielectric elastomer (DE) membranes.
-Among these approaches, deep reinforcement learning (DRL) stands out for its remarkable performance across diverse and complex tasks, offering a universal control solution.
-An increasing number of researchers have leveraged DRL algorithms to address control challenges in artificial muscle-driven robots, spanning applications ranging from dexterous grasping of robotic hands[20] to agile motion control of soft robotic manipulators [21, 22], and various others [23, 24, 25, 26].
-In most of existing studies [20, 22, 24, 23, 26], the standard procedure of applying DRL control in the targeted robotic systems typically follows the schematic illustrated in Fig.
+Over the past decade, robots driven by artificial muscles have achieved significant advancements, demonstrating impressive performance across diverse domains, including but not limited to biomimetic robots[2, 3], robot manipulators[4, 5], soft robots[6, 7], medical robot[8], origami robots[9], and marine robots[10].
+Previous research efforts have primarily centered around structural design and open-loop motion performance evaluation[2, 3, 4, 5, 6, 7, 8, 9, 10].
+Despite their potential, achieving accurate control of string-type artificial muscle-driven robots remains challenging due to their highly nonlinear and complex dynamics.
+Traditional control strategies, such as proportional-integral-derivative (PID) control[11, 12, 13], energy-based nonlinear control[14], and sliding mode control[15, 16] have been investigated to mitigate these challenges.
+Although these approaches have demonstrated success in specific tasks, a comprehensive, unified control strategy for string-type artificial muscle-driven robots is still lacking.
+As deep reinforcement learning (DRL) rapidly develops, learning-based controllers[17, 18, 19, 20, 21] have emerged as a promising alternative, offering the potential to overcome the limitations of conventional methods.
+The standard procedure of applying DRL in the targeted robotic systems typically follows the schematic illustrated in Fig.
 1.
-Initially, it is crucial to identify the relevant parameters of the artificial muscles and formulate the system dynamics model to establish a simulation environment.
-Subsequently, the DRL algorithms undergo training within simulation to develop a control policy.
-Finally, the acquired control policy is deployed on the robotic platform to accomplish various control objectives.
-Despite these advancements, challenges persist in the application of DRL control in the targeted robotic systems.
+Initially, identifying the relevant parameters and establishing a simulation environment enable a safe and efficient training process.
+The DRL algorithms are then trained within this simulation environment to develop an effective control policy.
+Finally, the learned control policy is transferred to the physical robotic platform to achieve the desired control objectives.
 Figure 1: Schematic of standard procedure of applying DRL control in artificial muscle-driven robotic systems.
-Three main steps include 1) dynamic modeling with system identification to establish simulation environment; 2) training DRL agent within the environment to develop a control policy; and 3) deploying the trained policy onto real-world robotic systems for testing.
-I-1 Muscle Sim-to-real Gap
-For the robotic systems driven by artificial muscles, modeling errors primarily come from the following three aspects.
-First, the dynamic models of artificial muscles themselves are approximate and inaccurate [27, 28].
-Second, inevitable errors occur in the system identification process of the muscles.
-Third, different muscle strings typically have different model parameters [29, 30], and conducting system identification for all muscle strings is time-consuming and labor-intensive.
-These factors jointly contribute to inaccurate modeling of artificial muscles, resulting in a significant sim-to-real gap for the entire robotic system.
-Consequently, control strategies that perform well in simulation may not seamlessly transfer to real-world applications [21, 22].
+However, several challenges persist in the applications of DRL to string-type artificial muscle-driven robots.
+One major issue is the sim-to-real gap, which arises from inaccuracies in dynamic models, system identification errors, and variability in muscle string parameters[22, 23].
+These factors often result in control strategies that demonstrate strong performance in simulation but struggle to transfer effectively to real-world applications.
+Another challenge is the low learning efficiency of DRL algorithms, which require substantial amounts of training data and iterations due to the complexity of neural networks and the trial and error learning process[24].
+Addressing these two challenges is crucial for advancing control technologies in artificial muscle-driven robots, which are increasingly employed in soft robotics and bio-inspired systems.
+This paper presents a DRL-based control strategy designed to enhance the data efficiency and real-world applicability of controllers for artificial muscle-driven robots.
+The proposed framework is evaluated on two robotic systems actuated by artificial muscle strings: a biomimetic robotic eye driven by super-coiled polymers (SCPs)[25, 26] and a parallel robotic wrist actuated by twisted-coiled actuators (TCAs)[27, 28].
+Specifically, we introduce a novel augmentation method for state-based data, which transforms a single transition data tuple into several, and we leverage an easy-to-implement PID controller to generate trajectories for early-stage training of DRL agents.
+Compared to the baseline algorithm (i.e., Soft Actor Critic [29]), the required training data is reduced by approximately $60\%$ in the targeted robotic systems.
+Additionally, we leverage a zero-shot sim-to-real transfer technique on artificial muscle dynamics.
+The experimental results demonstrate a significant improvement in real-world control performance with the averaged error reduced by $36.3\%$ through the application of this method.
+The remainder of this paper is organized as follows.
+Section II provides a review of the relevant literature.
+Section III introduces the basic concepts of applying reinforcement learning to the robotic systems studied.
+Section IV offers a concise overview of the robotic platforms utilized in this study.
+In Section V, we present a detailed description of our proposed efficient learning control framework.
+Section VI and Section VII discuss the experimental results, including training outcomes in simulation and control performances in real-world scenarios, respectively.
+Finally, concluding remarks and current limitations are discussed in Section VIII.
+II Related Work
+The control of artificial muscle-driven robotic systems using learning-based methods has garnered significant attention in recent years[30, 31, 20, 19, 32, 21, 18, 33].
+Among these approaches, DRL has demonstrated effectiveness in handling diverse and complex tasks, positioning it as a promising control solution for a wide range of applications.
+An increasing number of researchers have leveraged DRL algorithms to address control challenges in artificial muscle-driven robots, spanning applications ranging from dexterous grasping of robotic hands[17] to agile motion control of soft robotic manipulators[33, 18], and other related applications[32, 21].
+Many of these studies have noted that the low sample efficiency often leads to prolonged training time[20, 21], but few have focused specifically on improving training efficiency.
+Several existing state-of-the-art (SOTA) reinforcement learning (RL) methods may be applicable[34, 35, 36, 37], but they have not been evaluated within the context of the robotic systems we are addressing.
+The unique nonlinear and slow response characteristics of artificial muscle may present challenges that are not adequately addressed by these existing methods.
+Our previous work[19] explored low learning efficiency problem but was limited to simulation-based experiments.
+Due to the sim-to-real gap, control strategies that perform well in simulation may not seamlessly transfer to real-world applications[33, 18].
 Despite being labor-intensive, some researchers[31, 32] have attempted to directly train DRL algorithms on real robots.
-However, given the fragile nature of string-type artificial muscles [1], this approach lacks universality and safety guarantees.
-Hence, addressing the muscle sim-to-real gap for DRL-based controllers is imperative.
-I-2 Low Learning Efficiency
-In reinforcement learning, an agent continually interacts with the environment to acquire a good policy through learning from the interaction data [31].
-Due to the complexity and high-dimensional nature of deep neural networks, DRL algorithms usually require a large amount of training data and iterations for convergence towards the desired policy.
-While training in simulation remains feasible, many studies have noted that the low sample efficiency often leads to significantly prolonged training times [23, 26, 32].
-Moreover, the complex dynamic models of muscle-driven robots exacerbate this issue.
-Thus, enhancing training data efficiency and accelerating convergence speed are of great importance.
-We observe a scarcity of research addressing the aforementioned issues within the domain of artificial muscle-driven robots.
-Our prior research [24] aimed to tackle the challenge of low learning efficiency.
-However, its experimental findings were solely derived from simulations, without addressing the crucial muscle sim-to-real gap.
-This paper presents a DRL-based control strategy with improved data efficiency and transferability to real-world applications.
+However, the artificial muscles considered in this paper are all thermally actuated, which inherently leads to energy inefficiency and limited cycling rates[38].
+Continuous rapid transitions between heating and cooling often cause irreversible damage to the muscles, making direct training on real robots unsafe.
 It is worth noting that in recent years, there have been some efforts to bridge the sim-to-real gap of artificial muscles.
-Many researchers [27, 28, 33, 34, 35] have proposed more accurate yet significantly more complex dynamic models.
-In this paper, however, we seek to directly improve the performance of DRL-based controllers by employing a straightforward sim-to-real transfer technique, thereby bypassing the complicated process of mathematical modeling and parameter identification for artificial muscle strings.
-We evaluate our learning control frame work on two robotic systems powered by artificial muscle strings, including a biomimetic robotic eye driven by super-coiled polymers (SCPs) [36, 29], and a parallel robotic wrist actuated by twisted-coiled actuators (TCAs) [30, 37].
-The main contributions of this paper are as follows.
-To the best of our knowledge, this paper proposes the first efficient learning control framework designed for string-type artificial muscle-driven robotic systems.
-Specifically, we introduce a novel augmentation method for state-based data, which transforms a single transition data into several.
-We leverage an easy-to-implement PID controller to generate trajectories for early-stage training of DRL agents.
-The training results in simulation show that our proposed methods improve the data efficiency of the DRL algorithm in targeted robotic systems.
-Additionally, we leverage a zero-shot sim-to-real transfer technique on artificial muscle dynamics.
-The experimental results demonstrate a great improvement in real-world control performance through the application of this method in muscle-driven robotic systems.
-We believe that the proposed learning control framework can be utilized to address a variety of control tasks across similar robotic systems.
-The remainder of this paper is organized as follows.
-Section II presents basic information about leveraging reinforcement learning within the robotic systems of interest.
-Section III offers a concise overview of the robotic platforms utilized in this study.
-Section IV presents our proposed efficient learning control framework in detail.
-Subsequently, Section V presents the experimental results of the proposed method, encompassing training outcomes in simulation and control performances in real-world scenarios.
-Finally, concluding remarks and current limitations are provided in Section VI.
-II Preliminaries
-In this section, we model the control problem of targeted robotic systems as a reinforcement learning problem by Markov Decision Process (MDP), which is a mathematical framework that describes the interaction between the DRL agent and the environment [31].
-The MDP in artificial muscle-driven robots consists of the following tuple, i.e.,
-$(\mathcal{S},\mathcal{A},\mathcal{R},f,\gamma).$ (1)
+Many researchers[22, 23, 39] have proposed more accurate yet significantly more complex dynamic models.
+In this paper, however, we aim to directly improve the performance of DRL-based controllers by employing a sim-to-real transfer technique, thereby reducing the need for complicated mathematical modeling and precise parameter identification for artificial muscles.
+III Preliminaries
+In this section, we model the control problem of targeted robotic systems as a reinforcement learning problem by Markov Decision Process (MDP), which consists of the tuple $(\mathcal{S},\mathcal{A},\mathcal{R},f,\gamma)$.
   • $\mathcal{S}$ is the state space.
-At each time step $t$, the state vector $s_{t}\in\mathcal{S}$ generally includes the current state variables of the robotic system and the target positions and/or orientations.
-In this paper, the motion state variables of the robotic systems are collectively defined as vector $\bm{x}$, which contains the Euler angles, angular velocities, and other relevant motion states.
+The state vector $s_{t}\in\mathcal{S}$ generally includes the current state variables of the robotic system and the target positions and/or orientations.
+In this paper, the motion state variables of the robotic systems are collectively defined as vector $\bm{x}$.
 The system output vector, containing the variables to be controlled, is denoted by $\bm{y}$.
 Correspondingly, the target system output is denoted by $\bm{y}^{*}$.
-In this way, the state vector $s_{t}$ is represented as
-    $s_{t}=(\bm{x}_{t},\bm{y}^{*}).$ (2)
+In this way, the state vector $s_{t}$ is represented as $s_{t}=(\bm{x}_{t},\bm{y}^{*})$.
   • $\mathcal{A}$ is the action space, where each action $a_{t}\in\mathcal{A}$ is generated by the DRL agent based on the current state $s_{t}$ and its own policy.
-The action vector typically consists of the actuating voltage applied to each muscle or the voltagel difference between a pair of antagonistic muscles.
-  • $\mathcal{R}$ is the reward function that maps the current state $s_{t}$ and action $a_{t}$ to an immediate reward $r_{t}$, i.e.,
-    $r_{t}=\mathcal{R}(s_{t},a_{t}).$ (3)
+  • $\mathcal{R}$ is the reward function that maps the current state $s_{t}$ and action $a_{t}$ to an immediate reward $r_{t}=\mathcal{R}(s_{t},a_{t})$.
 The design of reward function plays a vital role in the learning process of the DRL agents, which learn the optimal policy under the guidance thereby.
-In this paper, we develop a generalized quadratic-form reward function design for the robotic systems of interest:
-    $r_{t}=-(e^{T}_{t}Q_{e}e_{t}+a^{T}_{t}R_{a}a_{t})+r_{\text{bonus}}$ (4)
+In this paper, we develop a generalized quadratic-form reward function designed for the robotic systems of interest:
+    $r_{t}=-(e^{\top}_{t}Q_{e}e_{t}+a^{\top}_{t}R_{a}a_{t})+r_{\text{bonus}},$ (1)
     where $e_{t}=|\bm{y}^{*}-\bm{y}_{t}|$ is the system tracking error, and $Q_{e}$ and $R_{a}$ represent the output and control weight matrices, respectively.
-A bonus reward $r_{\text{bonus}}$ is given when the system tracking error achieves a sufficiently small threshold.
-We add this bonus reward because, for the robotic systems discussed in this paper, maintaining at the target state is generally more challenging than reaching it.
-  • $f$ represents the state transition function, which describes the likelihood of a transition from state $s$ to $s^{\prime}$ when an action $a$ is taken.
-  • $\gamma\in[0,1]$ is the discount factor, which serves to balance the influence of future rewards on the current state.
-In all experiments conducted in this paper, we set $\gamma$ to a constant value of $\gamma=0.99$.
-III Experimental Platforms
-This section gives a concise overview of the testing platforms utilized in this paper, including an SCP-driven robotic eye (see Fig.
-2) and a TCA-driven parallel robotic wrist (see Fig.
-3).
-We briefly describe the hardware design, state space, action space, and reward function for each system.
+A bonus reward $r_{\text{bonus}}$ is given when the system tracking error falls below a sufficiently small threshold.
+  • $f$ represents the state transition function and $\gamma$ is the discount factor, which is set to $0.99$ in our experiments.
+Figure 2: Illustrations of SCP and TCA muscles.
+(a) Crimp fixtures for the SCP muscle ends (left), and a close-up shot of the fabricated SCP muscle (right); and (b) conductive ring with SMA wire wrapped around for the TCA muscle ends (left), and a close-up shot of the fabricated TCA muscle (right).
+IV Experimental Platforms
+This section presents an overview of the actuators and robotic platforms used in this paper, including SCP and TCA artificial muscles (Fig.
+2), a 2-DOF robotic eye (Fig.
+3) and a parallel robotic wrist (Fig.
+4).
+For each system, we briefly describe the hardware design, state space, action space, and reward function.
+IV-A Artificial Muscles
+Numerous fabrication methods for SCP and TCA muscles have been investigated in the literature[25, 26, 27, 28].
+Specifically, for our robotic eye design, the SCP actuator is fabricated by sequentially twisting, coiling, and double-backing using a 145 cm long silver-plated conductive nylon thread (Statex’s Shieldex 117/17 dtex 2 ply HC+B).
+As illustrated in Fig.
+2(a), the ends of the SCP muscles are crimped using fold-over silver/copper bead crimps.
+They are connected to N-Channel enhancement MOSFETs (FQP30N06L) which provide electric power to the muscles.
+Application of voltage induces Joule heating on the actuator which causes the coils to expand laterally, thus generating contracting force.
+Regarding our TCA muscles, the Shape Memory Alloy (SMA) wires (0.18 mm diameter) serve as the conductive threads and skeletons.
+Highly elastic spandex fibers (creora spandex, power fit, 840 denier) are tightly wound around the SMAs, and the resulting composite fibers are then twisted into a coil configuration.
+As shown in Fig.
+2(b), both ends of the SMA wire are wrapped around conductive rings, which are also connected to MOSFETs.
+When voltage potential is applied to the rings, the skeleton generates Joule heat and the surrounding spandex fibers will contract in longitudinal direction, which leads to the untwisting of the compound twisted fiber structure.
+The TCA’s coiled configuration converts this untwisting motion into linear contraction.
+Additionally, the SMA wire not only provides Joule heating, but also contributes additional contraction force owing to its shape memory effect.
 Figure 3: The 2-DOF robotic eye prototype driven by two pairs of SCP muscles.
-III-A Robotic Eye
+IV-B Robotic Eye
 The high occurrence of unsuccessful corrective surgeries for ocular motor disorders underscores the need for a muscle-driven robotic eye platform to assist ophthalmologists in gaining deeper insights into the biomechanics of human eye movement.
-In our previous work [36], we proposed a prototype of a 2-DOF robotic eye, and derived the geometric and dynamic models specific to this prototype.
+In our previous work[25], we proposed a prototype of a 2-DOF robotic eye and derived the geometric and dynamic models specific to this prototype.
 As illustrated in Fig.
-2, the hemispherical eyeball is driven by two pairs of SCP muscle actuators, accommodating a Raspberry Pi Mini Camera (RPiMC).
+3, the hemispherical eyeball is driven by two pairs of SCP muscle actuators, accommodating a Raspberry Pi Mini Camera (RPiMC).
 The primary control objective is to precisely and rapidly adjust the eyeball’s orientation towards the desired target angles by regulating the SCP muscles, namely foveation control in the eye movement study.
-The eyeball’s center of rotation, denoted as $\bm{O}\in\mathbb{R}^{3}$, remains fixed at the origin within the reference frame $O_{xyz}$.
-Here, the axes $x$, $y$, and $z$ correspond to the transverse, sagittal, and vertical axes of the eyeball, respectively.
+The eyeball’s center of rotation, denoted as $\bm{O}\in\mathbb{R}^{3}$, remains fixed at the origin within the reference frame $O_{\mathrm{xyz}}$.
+Here, the axes $\mathrm{x}$, $\mathrm{y}$, and $\mathrm{z}$ correspond to the transverse, sagittal, and vertical axes of the eyeball, respectively.
 Rotations around these axes are expressed as pitch $\theta$, roll $\phi$, and yaw $\psi$.
 The robotic eye system’s state variables are represented as
-$\bm{x}=[\theta,\dot{\theta},\psi,\dot{\psi},\Delta T_{1},\Delta T_{2},\Delta T_{3},\Delta T_{4}]$ (5)
+$\bm{x}=[\theta,\dot{\theta},\psi,\dot{\psi},\Delta T_{1},\Delta T_{2},\Delta T_{3},\Delta T_{4}]^{\top},$ (2)
 where $\theta,\dot{\theta},\psi,\dot{\psi},$ represent the pitch, yaw angles, and their respective angular velocities for control.
-Additionally, $\Delta T_{i}$ represents the temperature difference between the ambient temperature $T_{0}$ and the current temperature $T_{i}$ of muscle $m_{i}$.
-Figure 3: The parallel robotic wrist prototype actuated by three TCAs.
-Focusing primarily on the robotic eye’s orientation, the system’s output vector is denoted as $\bm{y}=[\psi,\dot{\psi},\theta,\dot{\theta}]$.
+$\Delta T_{i}$ represents the temperature difference between the ambient temperature $T_{0}$ and the current temperature $T_{i}$ of muscle $m_{i}$.
+Figure 4: The parallel robotic wrist prototype actuated by three TCA muscles.
+Focusing primarily on the robotic eye’s orientation, the system’s output vector is denoted as $\bm{y}=[\psi,\dot{\psi},\theta,\dot{\theta}]^{\top}$.
 Due to the difficulty and poor accuracy of measuring the temperature of each muscle in real time, the state vector of the DRL agent is defined by
-$s_{t}=[\theta_{t},\dot{\theta}_{t},\psi_{t},\dot{\psi}_{t},\theta^{*},\psi^{*}].$ (6)
+$s_{t}=[\theta_{t},\dot{\theta}_{t},\psi_{t},\dot{\psi}_{t},\theta^{*},\psi^{*}]^{\top}.$ (3)
 Here, $\theta^{*}$ and $\psi^{*}$ represent the target angles.
@@
 Given the antagonistic configuration of the muscle pairs, the control input is defined by a constrained action vector, i.e.,
-$a_{t}=[$ $-V_{1}+V_{2},-V_{3}+V_{4}]$ (7)
-s.t. $V_{1}\cdot V_{2}=V_{3}\cdot V_{4}=0$
-$0\leq V_{i}\leq 10$
+$a_{t}=[$ $-V_{1}+V_{2},-V_{3}+V_{4}]^{\top}$ (4)
+[…]
 The constraint requires that only one muscle is activated at each time step in the pairs $(m_{1},m_{2})$ and $(m_{3},m_{4})$.
-The reward function follows the structure of (4), where the output and control weight matrices are set to $Q_{e}=\text{diag}\{0.05,0.25,0.05,0.25\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively.
+The reward function follows the structure of (1), where the output and control weight matrices were set to $Q_{e}=\text{diag}\{0.05,0.25,0.05,0.25\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively.
+They were chosen based on empirical tuning to balance the trade-off between output error and control effort.
 The bonus reward is granted when the tracking error is sufficiently small, given by
@@
 Here, $\mathds{1}\{\cdot\}$ denotes the binary indicator function that yields a value of $1$ if the given condition holds true, and 00 otherwise.
+IV-C Robotic Wrist
+Research into artificial muscle-driven robotic wrist joints is of great importance to several domains, including the development of dexterous robotic systems with human-like capabilities, advancements in rehabilitation and prosthetics technology, and the enhancement of robotic manipulation versatility across various fields.
+We designed a parallel robotic wrist prototype driven by three equally spaced compound TCAs[27] as illustrated in Fig.
+4.
+The base plate remains stationary, while the upper plate, also known as the end effector, moves via three muscle strings.
+Linkages between the two plates are constructed to support a two degree-of-freedom motion of the upper plate[40].
+Given this configuration, we developed a simplified dynamics model under the assumption of a fixed rotational center for the upper plate.
+Similar to the foveation control of the robotic eye, the control objective entails precise regulation of the moving upper plate’s orientation.
+This necessitates adjusting the applied actuating voltages across the three muscle strings to drive the upper plate from its initial equilibrium state to predefined target angles.
+The reference frame $O_{\mathrm{xyz}}$ is fixed on the moving upper plate, with the $\mathrm{z}$-axis perpendicular to the upper plate and pointing outward.
+$\theta,\phi,$ and $\psi$ represent the pitch, roll, and yaw angles, respectively.
+Similar to the robotic eye, the system state variables are defined as
+$\bm{x}=[\theta,\dot{\theta},\phi,\dot{\phi},\Delta T_{1},\Delta T_{2},\Delta T_{3}]^{\top},$ (6)
+where $\dot{\theta}$ and $\dot{\phi}$ represent the corresponding angular velocities, and $\Delta T_{i}$ is the temperature difference of muscle $m_{i}$ with respect to the ambient environment, with $i\in\{1,2,3\}$.
+The system output vector is represented as $\bm{y}=[\theta,\dot{\theta},\phi,\dot{\phi}]^{\top}$.
+Likewise, the muscle’s temperature is not used as an input for the DRL agent, and thus, the state vector is defined as
+$s_{t}=[\theta_{t},\dot{\theta}_{t},\phi_{t},\dot{\phi}_{t},\theta^{*},\phi^{*}]^{\top},$ (7)
+where $\theta^{*}$ and $\phi^{*}$ denote the target angles, with the respective Euler angles and angular velocities measured using the inertial measurement unit (IMU) positioned at the center of the upper plate.
+Additionally, the input action vector $a_{t}$ of the system is defined as:
+[…]
+The output and control weight matrices of the reward function are set to $Q_{e}=\text{diag}\{0.05,0.2,0.05,0.2\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively.
+When the tracking error is sufficiently small, a bonus reward is granted, i.e.,
+$r_{\text{bonus}}=2\times\left(\mathds{1}\{|\theta-\theta^{*}|<0.5\}+\mathds{1}\{|\phi-\phi^{*}|<0.5\}\right).$ (9)
+V Learning Control Framework
 Figure 5: Overview of the proposed efficient learning control framework, including a two-stage training process in simulation and real-world application.
@@
 Stage 2 integrates the baseline SAC learning algorithm with state-based data augmentation and randomization of muscle dynamics for expediting training convergence and bridging sim-to-real gap, respectively.
-III-B Robotic Wrist
-Research into artificial muscle-driven robotic wrist joints holds considerable importance in several domains, including the development of dexterous robotic systems with human-like capabilities, advancements in rehabilitation and prosthetics technology, and the enhancement of robotic manipulation versatility across various fields.
-We designed a parallel robotic wrist prototype driven by three equally spaced compound TCAs [30] as illustrated in Fig.
-3.
-The base plate remains stationary, while the upper plate, also known as the end effector, moves via three muscle strings.
-Linkages between the two plates are constructed to support two degree-of-freedom motion of the upper plate [38].
-Given this configuration, we developed a simplified dynamics model under the assumption of a fixed rotational center for the upper plate.
-Similar to the foveation control of the robotic eye, the control objective entails precise regulation of the moving upper plate’s orientation.
-This necessitates adjusting the applied actuating voltages across the three muscle strings to drive the upper plate from its initial equilibrium state to predefined target angles.
-The reference frame $O_{xyz}$ is fixed on the moving upper plate, with the $z$-axis perpendicular to the upper plate and pointing outward.
-$\theta,\phi,$ and $\psi$ represent the pitch, roll, and yaw angles, respectively.
-Similar to the robotic eye, the system state variables are defined as
-$\bm{x}=[\theta,\dot{\theta},\phi,\dot{\phi},\Delta T_{1},\Delta T_{2},\Delta T_{3}]$ (9)
-where $\dot{\theta}$ and $\dot{\phi}$ represent the corresponding angular velocities, and $\Delta T_{i}$ is the temperature difference of muscle $m_{i}$ with respect to the ambient environment, with $i\in\{1,2,3\}$.
-The system output vector is represented as $\bm{y}=[\theta,\dot{\theta},\phi,\dot{\phi}]$.
-Likewise, the muscle’s temperature is not used as an input for the DRL agent, and thus, the state vector is defined as
-$s_{t}=[\theta_{t},\dot{\theta}_{t},\phi_{t},\dot{\phi}_{t},\theta^{*},\phi^{*}].$ (10)
-Here, $\theta^{*}$ and $\phi^{*}$ denote the target angles, with the respective Euler angles and angular velocities measured using the inertial measurement unit (IMU) positioned at the center of the upper plate.
-Additionally, the input action vector $a_{t}$ of the system is defined as
-$a_{t}=[V_{1},V_{2},V_{3}]$ (11)
-where $V_{i}$ represents the actuating voltage applied to muscle $m_{i}$, constrained by $0\leq V_{i}\leq 10$.
-The output and control weight matrices of the reward funtion are set to $Q_{e}=\text{diag}\{0.05,0.2,0.05,0.2\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively.
-When the tracking error is sufficiently small, a bonus reward is granted, i.e.,
-$r_{\text{bonus}}=2\times\left(\mathds{1}\{|\theta-\theta^{*}|<0.5\}+\mathds{1}\{|\phi-\phi^{*}|<0.5\}\right).$ (12)
-IV Learning Control Framework
 This section provides a comprehensive overview of the proposed learning control framework, as illustrated in Fig.
-4.
-The soft actor-critic (SAC) serves as the baseline DRL algorithm.
-To enhance data efficiency, we employ a two-stage training strategy.
+5.
+The soft actor-critic (SAC) algorithm serves as the baseline DRL method.
+To improve data efficiency, we employ a two-stage training strategy.
 Initially, a standard PID controller is initialized to generate imperfect trajectories for the DRL agent.
 Subsequently, the agent interacts directly with the simulation environment to update its policy.
-All transition data observed from the environment undergoes an state-based augmentation strategy before being stored into the replay buffer.
-To address the sim-to-real gap, we introduce a randomization method on the artificial muscle dynamics at the beginning of each training episode.
-Eventually, the trained actors (control policies) are able to be directly applied in real-world robotic systems to achieve desired control performance.
-IV-A Soft Actor-Critic (SAC)
-SAC, a state-of-the-art deep reinforcement learning algorithm, finds extensive application across various tasks [39]. Traditional DRL algorithms primarily maximize cumulative rewards, i.e.,
-$J(\pi)=\sum_{t=0}^{T}\mathbb{E}_{(s_{t},a_{t})\sim\rho_{\pi}}[\mathcal{R}(s_{t},a_{t})]$ (13)
-where $\pi$ represents the policy. However, SAC adopts an entropy-maximizing approach, with its learning objective formulated as
-$J(\pi)=\sum_{t=0}^{T}\mathbb{E}_{(s_{t},a_{t})\sim\rho_{\pi}}[\mathcal{R}(s_{t},a_{t})+\alpha\mathcal{H}(\pi(\cdot|s_{t}))].$ (14)
-where $\mathcal{H}(\pi(\cdot|s_{t}))$ represents an entropy of the policy $\pi$, and $\alpha$ serves as a temperature parameter.
-This approach enhances the robustness of DRL and accelerates its training process [31].
-Figure 5: Schematic diagram of the network architecture.
-The actor (left) and the critic (right) networks share the same structure but have different parameters.
-Each is composed of a fully-connected layer with ReLU activation as the input layer, a recurrent GRU [40] unit, and a fully-connected output layer.
-In addition, SAC is an off-policy algorithm that leverages the actor-critic architecture.
-The neural network structures utilized in this paper are illustrated in Fig.
-5.
-To process the sequential information crucial for implementing dynamics randomization and handling partial observability, we employ a Gated Recurrent Unit (GRU) [40].
-We select GRU over over commonly-used Long Short-Term Memory (LSTM) due to its efficiency and ease of training [40].
-Noise is added to the observation input to the actor.
-The hyperparameters of the SAC algorithm are detailed in Table I and remain consistent across all ablation studies conducted in this paper.
-TABLE I: SAC Hyper-parameters
-Parameter Value
-Learning rate of actor $1\times 10^{-4}$
-Learning rate of critic $1\times 10^{-4}$
-Replay buffer size $1\times 10^{5}$
-Mini-batch size $20$
-Entropy target $-\text{dim}(\mathcal{A})$ (e.g., $-3$ for robotic wrist)
-Soft update factor $\tau$ $0.005$
-Discount factor $\gamma$ $0.99$
-Hidden layer size of GRU $256$
-Gradient steps $k$ $40$
-Target networks update interval $1$
-IV-B Randomization of Muscle Dynamics
+All transition data collected from the environment undergoes a state-based augmentation strategy before being stored into the replay buffer.
+To mitigate the sim-to-real gap, we introduce a randomization method on the artificial muscle dynamics at the beginning of each training episode.
+Eventually, the trained actors (control policies) can be directly implemented in real-world robotic systems to achieve the desired control performance.
+V-A Soft Actor-Critic
+Soft Actor-Critic (SAC) [29], a state-of-the-art DRL algorithm, is widely adopted as a strong baseline for continuous control tasks in robotics due to its robustness, sample efficiency, and model-free nature.
+Unlike traditional DRL algorithms, SAC incorporates an entropy-maximizing approach, with its learning objective defined as:
+$J(\pi)=\sum_{t=0}^{T}\mathbb{E}_{(s_{t},a_{t})\sim\rho_{\pi}}\left[\mathcal{R}(s_{t},a_{t})+\alpha\mathcal{H}(\pi(\cdot|s_{t}))\right],$ (10)
+where $\mathcal{H}(\pi(\cdot|s_{t}))$ represents the entropy of the policy $\pi$, and $\alpha$ serves as a temperature parameter.
+This entropy-regularized objective improves the robustness of DRL and accelerates its training [24].
+To process sequential information critical for addressing partial observability issue and implementing dynamics randomization, we integrate a Gated Recurrent Unit (GRU) [41] as the hidden layer in the actor-critic architecture.
+V-B Randomization of Muscle Dynamics
+TABLE I: Dynamics parameters of randomization and their respective ranges in the robotic eye
+Parameter Value Scailing factor Additive term
+[…]
+[…]
+$c$ $(\text{N}/^{\circ}\text{C})$ $0.0055$ uniform$(0.85,1.15)$
+$C_{\text{th}}$ $(\text{Ws}/^{\circ}\text{C})$ $0.28$ uniform$(0.8,1.2)$
+$\lambda$ $(\text{W}/^{\circ}\text{C})$ $0.094$ uniform$(0.85,1.15)$
+$R$ $(\Omega)$ $20.0$ uniform$(0.9,1.1)$
+$\theta,\psi$ noise $(^{\circ})$ normal$(0,0.1)$
+$\dot{\theta},\dot{\psi}$ noise $(^{\circ}/\text{s})$ normal$(0,0.05)$
+The dynamic model of the artificial muscles considered in this paper is based on the literature[26], where the muscle’s elasticity behavior is linearized to establish a dynamical system comprising a mechanical model characterized by stiffness and damping, given by
+$F=k(x-x_{0})+b\dot{x}+c(T(t)-T_{\text{amb}}),$ (11)
+where $x$ and $x_{0}$ denote the loaded length and unloaded length of the actuator, respectively, while $k$, $b$, and $c$ denote the mean stiffness, damping, and temperature effect coefficients of the actuator, respectively.
+$T(t)$ denotes the current actuator temperature, while $T_{\text{amb}}$ stands for the ambient temperature of the environment.
+The thermoelectric model of the actuator is expressed as:
+$C_{\text{th}}\frac{dT(t)}{dt}=P(t)-\lambda(T(t)-T_{\text{amb}}),$ (12)
+where $C_{\text{th}}$ represents the thermal mass of the actuator (measured in $\text{Ws}/^{\circ}\text{C}$), $P(t)$ indicates the heat applied to the actuator, $\lambda$ is the absolute thermal conductivity of the actuator with respect to the ambient environment (in $\text{W}/^{\circ}\text{C}$).
+The applied power, $P(t)$, is controlled through the voltage applied to the muscle, expressed as $P(t)=V(t)^{2}/R$.
+TABLE II: Dynamics parameters of randomization and their respective ranges in the robotic wrist
+Parameter Value Scailing factor Additive term
+[…]
+[…]
+$c$ $(\text{N}/^{\circ}\text{C})$ $0.0707$ uniform$(0.85,1.15)$
+$C_{\text{th}}$ $(\text{Ws}/^{\circ}\text{C})$ $3.06$ uniform$(0.8,1.2)$
+$\lambda$ $(\text{W}/^{\circ}\text{C})$ $0.1189$ uniform$(0.85,1.15)$
+$R$ $(\Omega)$ $10.0$ uniform$(0.9,1.1)$
+$\theta,\phi$ noise $(^{\circ})$ normal$(0,0.1)$
+$\dot{\theta},\dot{\phi}$ noise $(^{\circ}/\text{s})$ normal$(0,0.05)$
 The core idea behind dynamics randomization is to introduce randomness into specific dynamic parameters within the simulator during training.
 This approach allows the DRL agent to learn a policy adaptable to various dynamics conditions, including those encountered in real-world experiments.
-Widely adopted in diverse robotic applications [41, 42, 43], this method serves as inspiration for our work.
-Specifically, we randomize several critical parameters within the simulated dynamics environment.
-Among the robotic systems discussed in this paper, the dynamic parameter set of the muscles, denoted by $\mu$, holds pivotal importance.
-Therefore, we establish a distribution $\rho_{\mu}$ to encompass the variability of muscle dynamic parameters for dynamics randomization purposes.
-Additionally, we incorporate Gaussian noise into the observation inputs of the policy to emulate the noise commonly present in real-world sensor data.
-Regarding the SCP muscles driving the robotic eye and the TCA muscles driving the parallel robotic wrist, their thermomechanical models are approximately as [29]
-$F=k(x-x_{0})+b\dot{x}+c(T(t)-T_{\text{amb}}).$ (15)
-Here, $x$ and $x_{0}$ represent the loaded length and unloaded length of the actuator, respectively.
-$k$, $b$, and $c$ denote the mean stiffness, damping, and temperature effect of the actuator, respectively.
-$T(t)$ denotes the current temperature, while $T_{\text{amb}}$ stands for the ambient temperature of the environment.
-The thermoelectric model of the actuator is expressed as
-$C_{\text{th}}\frac{dT(t)}{dt}=P(t)-\lambda(T(t)-T_{\text{amb}}).$ (16)
-Here, $C_{\text{th}}$ represents the thermal mass of the actuator, and $\lambda$ represents the absolute thermal conductivity of the actuator.
-$P(t)$ stands for the heat power applied to the actuator, calculated as $P(t)=V^{2}/R$, where $V$ is the voltage applied to the actuator and $R$ is the electrical resistance.
-In summary, the dynamic parameter set for randomization is
-$\mu:=\{k,b,c,C_{\text{th}},\lambda,R\}$ (17)
-TABLE II: Dynamics parameters of randomization and their respective range in the robotic eye
-Parameter Value Scailing factor Additive term
-$k$ $(N/cm)$ $0.25$ uniform$(0.8,1.2)$
-$b$ $(N\cdot s/cm)$ $0.01$ uniform$(0.9,1.1)$
-$c$ $(N/^{\circ}C)$ $0.0055$ uniform$(0.85,1.15)$
-$C_{\text{th}}$ $(Ws/^{\circ}C)$ $0.28$ uniform$(0.8,1.2)$
-$\lambda$ $(W/^{\circ}C)$ $0.094$ uniform$(0.85,1.15)$
-$R$ $(\Omega)$ $20.0$ uniform$(0.9,1.1)$
-$\theta,\psi$ noise $(^{\circ})$ normal$(0,0.1)$
-$\dot{\theta},\dot{\psi}$ noise $(^{\circ}/s)$ normal$(0,0.05)$
-Table II and III outline the distribution $\rho_{\mu}$ of SCP and TCA muscles, respectively, along with the observation noise added for the robotic eye and wrist systems.
+Specifically, we randomize several critical muscle parameters within the simulated dynamics environment.
+A distribution $\rho_{\mu}$ of the muscle parameter set $\mu$ is established to encompass the variability of muscle dynamic parameters for dynamics randomization purposes.
+This approach is expected to be particularly effective in our setup, as artificial muscles are known to exhibit significant variability due to factors such as material degradation, slight manufacturing differences, and environmental temperature changes.
+According to the muscle models (11) and (12), the dynamic parameter set $\mu$ for randomization is given by
+$\mu:=\left\{k,b,c,C_{\text{th}},\lambda,R\right\}.$ (13)
+Table I and II outline the distribution $\rho_{\mu}$ of SCP and TCA muscles, respectively, along with the observation noise added for the robotic eye and wrist systems.
 $\rho_{\mu}$ is centered on the mean values derived from system identification to align the data distribution in simulation more accurately with that in reality.
-The range of randomization for each parameter is determined based on the distribution observed across different real muscle strings, as outlined in previous work [29].
-During the training process, the physical parameters of the actuators are randomized within the scaling ranges (i.e., $\mu\sim\rho_{\mu}$) at the beginning of each training episode and remain constant throughout.
-Whereas, the observation noise varies at each time step and across different Euler angles.
-While this paper employs a simplified linear dynamic model for artificial muscles, this method is expected to be readily transferred to more complex muscle dynamics models.
-IV-C Bootstrap via PID
-On-line DRL agents mainly rely on random exploration [31].
-However, due to the slow response of artificial muscles, random exploration policies may struggle to explore the state space comprehensively and rapidly, resulting in low sample efficiency.
-To address this issue, we draw inspiration from previous work [44, 45] and boost the training process by providing demonstrations into initial replay buffer.
-While PID controllers demand extensive tuning for optimal control performance, their implementation is relatively straightforward, given that moderate performance suffices.
-In the case of muscle-driven robotic systems, PID policies typically suffice to approach the target state.
-Hence, they provide trajectories that explore the defined state space more comprehensively, serving as an advantageous starting point for the DRL learning process.
-Subsequently, the agent will interact with the environment and continue to refine its policy, thus eventually outperforming PID and achieving desired control performances.
-Before the DRL training process begins, we initialize a PID controller that offers a barely acceptable (and typically not excellent) control performance.
-Within the first $M$ episodes of training, the PID controller interacts with the environment and generates trajectories, which are then stored in the replay buffer along with the the rewards.
-Differing from traditional behavior cloning methods, we don’t utilize a supervised loss during demonstration learning.
-Consequently, the agent does not strictly follow the PID control policy but leverages the trajectories generated by PID solely to bootstrap the initial training.
-IV-D Augmentation of State Vector
-Data augmentation refers to a class of techniques aimed at enhancing the size and quality of training datasets, thereby facilitating the development of more robust deep learning models [46].
-The fundamental concept of data augmentation involves transforming limited data into diverse, functionally equivalent datasets.
+To strike the right balance between randomization and fidelity, we carefully tune the range and distribution of randomized parameters based on empirical observations of muscle behavior in real-world scenarios.
+Specifically, the range of each parameter is determined based on the distribution observed across different real muscle strings, as outlined in [26].
+This approach allows us to maintain a high degree of fidelity while simultaneously enabling the controller to generalize effectively to real-world uncertainties.
+Details on the impact of varying randomization ranges on experimental results are discussed in Section VII-C.
+During the training process, the physical parameters of the actuators are randomized within predefined scaling ranges (i.e., $\mu\sim\rho_{\mu}$) at the start of each training episode and remain constant throughout.
+Additionally, Gaussian noise is added to the policy’s observation inputs to simulate the noise typically present in real-world sensor data.
+This observation noise varies at each time step and across different state variables.
+Although this study employs a simplified dynamic model for artificial muscles, the proposed method can be extended to more complex muscle dynamics models to further enhance control performance.
+V-C Bootstrap via PID
+On-line DRL agents primarily rely on random exploration [24].
+However, due to the slow response characteristics of artificial muscles, random exploration policies often fail to explore the state space comprehensively and efficiently, leading to low sample efficiency.
+To address this limitation, we accelerate the early-stage training by incorporating demonstration data into the initial replay buffer.
+For muscle-driven robotic systems, PID controllers are typically sufficient to approach target states and deliver moderate control performance.
+As such, they can generate trajectories that explore the defined state space more comprehensively, providing a beneficial starting point for the DRL learning process.
+Specifically, we initialize a PID controller that offers a barely acceptable (and typically not excellent) control performance before DRL training begins.
+Within the first $M$ episodes of training, the PID controller interacts with the environment and generates trajectories, which are then stored in the replay buffer along with the rewards.
+Unlike traditional behavior cloning methods, we do not employ a supervised loss for demonstration learning, allowing the agent to avoid strictly mimicking the PID control policy.
+Subsequently, the agent continues interacting with the environment to refine its policy, eventually surpassing the PID controller and achieving superior control performance.
+V-D Augmentation of State Vector
+Data augmentation refers to a class of techniques aimed at increasing the size and improving the quality of training datasets, thereby facilitating the development of more robust deep learning models[42].
+The fundamental concept of data augmentation involves transforming limited data into diverse and functionally equivalent datasets.
 Given the labor-intensive and computationally expensive nature of collecting training data in robotic systems, we integrate data augmentation into the agent’s learning process.
-While prior research [47, 48, 49] has explored data augmentation in reinforcement learning, existing studies have primarily focused on pixel-based RL scenarios, where agents rely on image-based observations.
-In this paper, we present a straightforward and adaptable approach to state-based data augmentation, better suited and more effective for our targeted robotic systems.
-TABLE III: Dynamics parameters of randomization and their respective range in the robotic wrist
-Parameter Value Scailing factor Additive term
-$k$ $(N/mm)$ $0.21$ uniform$(0.8,1.2)$
-$b$ $(N\cdot s/cm)$ $0.63$ uniform$(0.9,1.1)$
-$c$ $(N/^{\circ}C)$ $0.0707$ uniform$(0.85,1.15)$
-$C_{\text{th}}$ $(Ws/^{\circ}C)$ $3.06$ uniform$(0.8,1.2)$
-$\lambda$ $(W/^{\circ}C)$ $0.1189$ uniform$(0.85,1.15)$
-$R$ $(\Omega)$ $10.0$ uniform$(0.9,1.1)$
-$\theta,\phi$ noise $(^{\circ})$ normal$(0,0.1)$
-$\dot{\theta},\dot{\phi}$ noise $(^{\circ}/s)$ normal$(0,0.05)$
-As commonly acknowledged, rewards serve as pivotal guides in the learning process of DRL agents.
-Therefore, our objective is to generate more data transitions with different rewards.
-By combining (2) and (3), the reward $r_{t}$ is then calculated as
-$r_{t}=\mathcal{R}(\bm{x}_{t},\bm{y}^{*},a_{t}).$ (18)
-Here, $\bm{x}_{t}$ and $a_{t}$ are directly related to system dynamics, whereas the target output vector $\bm{y}^{*}$ is manually set and changeable.
-Consequently, the proposed augmentation of state vector proceeds as follows.
-At each time step $t$, we fix the motion state vectors $\bm{x}_{t},\bm{x}_{t+1}$ and the action $a_{t}$, and vary the target vector $\bm{y}^{*}$.
-In this paper, we introduce a random additive term into $\bm{y}^{*}$ as follows
-$\bm{y}^{*}_{\text{new}}=\bm{y}^{*}\pm\delta\bm{Z}.$ (19)
-Here, $\bm{y}^{*}_{\text{new}}$ denotes the new target vector.
-$\delta$ acts as a scalar factor to control the scope of augmentation, which largely depends on the defined range.
-$\bm{Z}$ is a random vector sharing identical dimensions with $\bm{y}^{*}$, wherein the elements of $\bm{Z}$ range from 00 to $1$.
-Thus, we derive a pair of new states $s^{\prime}_{t}$ and $s^{\prime}_{t+1}$.
-After that, a new corresponding reward $r^{\prime}_{t}$ is calculated using (18).
-Following the above steps, the original transition tuple is transformed into a new one, i.e.,
-$(s_{t},a_{t},r_{t},s_{t+1})\xrightarrow{AUG}(s^{\prime}_{t},a_{t},r^{\prime}_{t},s^{\prime}_{t+1}).$ (20)
-Algorithm 1 Augmentation of State-based Data
-1:  Generate rollout $\tau=(s_{0},a_{0},r_{0},...,s_{T})$ from environment
-2:  for $i=1$ to $n$ do
-3:     for each $s_{t},a_{t},r_{t}$ in $\tau$ do
-4:        Create a new state vector $s_{t}\rightarrow s^{\prime}_{t}$
-5:        Calculate a new reward $r^{\prime}_{t}=\mathcal{R}(s^{\prime}_{t},a_{t})$
-6:     end for
-7:     Get a new rollout $\tau^{\prime}_{i}=(s^{\prime}_{0},a_{0},r^{\prime}_{0},...,s_{T})$
-8:  end for
-TABLE IV: Simulation Parameters of Robotic Systems
-Description Symbol Value
-Robotic Eye Total time duration of one episode $t_{\text{total}}$ $15$s
-Action time step $t_{a}$ $0.5$s
-Total number of training episodes $N$ $2000$
-Random or boost episodes $M$ $250$
-Absolute value of $\theta$ bound $\theta_{\text{max}}$ $15^{\circ}$
-Absolute value of $\psi$ bound $\psi_{\text{max}}$ $10^{\circ}$
-Robotic Wrist Total time duration of one episode $t_{\text{total}}$ $30$s
-Action time step $t_{a}$ $0.5$s
-Total number of training episodes $N$ $3500$
-Random or boost episodes $M$ $500$
-Absolute value of $\theta$ bound $\theta_{\text{max}}$ $15^{\circ}$
-Absolute value of $\phi$ bound $\phi_{\text{max}}$ $15^{\circ}$
+This paper proposes a novel and straightforward state-based data augmentation technique designed for our targeted robotic systems.
 Algorithm 1 SAC-BAR Learning Control
@@
 32:  end for
-The procedure of using augmentation for a state-based trajectory $\tau$ is provided in Algorithm 1.
-Each time the agent interacts with the environment and gets a new trajectory, this method creates $n$ different trajectories at a very low computational cost.
-We set $n=10$ in both robotic eye and wrist systems during training.
-In this paper, we denote the bootstrap method by the letter ’B’, augmentation method by the letter ’A’, and randomization of muscle dynamics by the letter ’R’.
-Algorithm 2 provides a detailed overview of the entire training procedure for the SAC algorithm with all three proposed improvements (i.e., SAC-BAR).
-V Experimental Results
-This section presents the experimental results of the proposed efficient learning control framework within two testing robotic platforms.
-We first present the training results of the algorithm in simulation environments, validating the effectiveness of the proposed bootstrap and augmentation methods in enhancing training data efficiency.
-Subsequently, we demonstrate the results of an ablation study conducted on the robotic wrist platform, verifying the effectiveness of dynamics randomization in improving the real-world control performance.
-Finally, experimental results on robotic eye are provided to demonstrate the generalizability of the proposed control framework across different robotic systems of interest.
-(a) Robotic Eye
-(b) Robotic Wrist
-Figure 6: Reward trajectories with respect to the training episode number in ablation study.
-Ours refers to the proposed SAC-BAR algorithm, where B, A, and R indicate bootstrap, augmentation, and randomization techniques, respectively.
-Performance is evaluated in simulation with random dynamics every $10$ training episodes.
-The shaded area indicates the standard deviation of the mean reward based on running each experiment with three different random seeds.
-Figure 7: Experimental trajectories of vanilla SAC algorithm on the robotic wrist under target angles ($\theta^{*}=-5^{\circ},\phi^{*}=7.5^{\circ}$) and ($\theta^{*}=2.5^{\circ},\phi^{*}=-7.5^{\circ}$), respectively.
-(a), (c) The system output angles $(\theta,\phi)$.
-(b), (d) The corresponding actions $a_{t}$ generated by the trained actor.
-Figure 8: Experimental trajectories of our proposed SAC-BAR algorithm on the robotic wrist under target angles ($\theta^{*}=-5^{\circ},\phi^{*}=7.5^{\circ}$) and ($\theta^{*}=2.5^{\circ},\phi^{*}=-7.5^{\circ}$), respectively.
-(a), (c) The system output angles $(\theta,\phi)$.
-(b), (d) The corresponding actions $a_{t}$ generated by the trained actor.
-V-A Training Results
-We established simulation environments for two robotic systems to train DRL agents and conduct preliminary tests.
-Table IV lists the simulation parameters for each robotic system.
-The overall duration of a single episode $t_{\text{total}}$ and the action time step $t_{a}$ depend on the rise time and response speed constraints inherent to the hardware of the systems.
-The values of $N$ and $M$ are empirical parameters contingent on the complexity of the systems.
-For simpler dynamics, smaller values for $N$ and $M$ facilitate the learning process for the DRL agent in acquiring an effective control policy.
-For algorithms not using the bootstrap method, random exploration is employed in the first $N$ episodes.
-Target angles are confined within the motion range of the robotic device.
-With the majority of training time spent on dynamics calculation, the time step for dynamics simulation is set at $0.01$s to balance simulation accuracy and computational efficiency.
+As commonly acknowledged, rewards serve as pivotal guides in the learning process of DRL agents.
+Therefore, our objective is to generate additional data transitions with different rewards.
+As described in Section III, the reward $r_{t}$ is calculated as:
+$r_{t}=\mathcal{R}(\bm{x}_{t},\bm{y}^{*},a_{t}).$ (14)
+Here, $\bm{x}_{t}$ and $a_{t}$ are directly related to system dynamics, whereas the target output vector $\bm{y}^{*}$ is manually set and changeable.
+Consequently, the proposed augmentation of state vectors proceeds as follows.
+At each time step $t$, we fix the motion state vectors $\bm{x}_{t},\bm{x}_{t+1}$ and the action $a_{t}$, and vary the target vector $\bm{y}^{*}$.
+In this paper, we introduce a random additive term into $\bm{y}^{*}$ as follows
+$\bm{y}^{*}_{\text{new}}=\bm{y}^{*}\pm\delta\bm{Z}.$ (15)
+Here, $\bm{y}^{*}_{\text{new}}$ denotes the new target vector.
+$\delta$ acts as a scalar factor to control the scope of augmentation, which largely depends on the defined range.
+$\bm{Z}$ is a random vector sharing identical dimensions with $\bm{y}^{*}$, wherein the elements of $\bm{Z}$ range from 00 to $1$.
+Thus, we derive a pair of new states $s^{\prime}_{t}$ and $s^{\prime}_{t+1}$.
+After that, a new corresponding reward $r^{\prime}_{t}$ is calculated using (14).
+Following the above steps, the original transition tuple is transformed into a new one, i.e.,
+$(s_{t},a_{t},r_{t},s_{t+1})\xrightarrow{AUG}(s^{\prime}_{t},a_{t},r^{\prime}_{t},s^{\prime}_{t+1}),$ (16)
+which creates different transition tuples at a negligible computational cost.
+In this paper, we denote the bootstrap method by the letter ’B’, augmentation method by the letter ’A’, and the randomization of muscle dynamics by the letter ’R’.
+Thus, our proposed SAC algorithm with all three improvements is denoted by SAC-BAR.
+The pseudo code of our overall SAC-BAR learning control framework is presented in Algorithm 1.
+Figure 6: Reward trajectories with respect to the training episode number.
+(a) Comparison results between our SAC-BAR and SOTA methods including REDQ, DroQ, RLPD, and X-QL.
+(b) Results of ablation study where ‘B’, ‘A’, and ‘R’ indicate bootstrap, augmentation, and randomization techniques, respectively.
+VI Simulation Experiments
+In this section, we present the training results from simulation environments, comparing our proposed SAC-BAR against several SOTA reinforcement learning algorithms in terms of sample efficiency on two robotic systems.
+Additionally, we perform ablation studies to evaluate the effectiveness of the proposed bootstrap and augmentation methods.
+VI-A Training Setup
+At the beginning of each training episode, a target state is randomly selected from a uniformly distributed range.
+For instance, in the robotic eye scenario, the target angles adhere to $\theta^{*}\sim\mathcal{U}[-15^{\circ},15^{\circ}]$ and $\psi^{*}\sim\mathcal{U}[-10^{\circ},10^{\circ}]$.
+The robotic agent strives to reach the target from the initial state throughout the episode.
+In this paper, simulation environments are all written in Python.
 Training and testing in simulation are conducted on a computer with an Intel Xeon Platinum 8383C CPU and an NVIDIA GeForce RTX 4090 GPU.
 The algorithms are implemented in PyTorch and trained using the Adam optimizer.
-In each training episode, a target state is randomly selected from a uniformly distributed range.
-For instance, in the robotic eye scenario, the target angles adhere to $\theta^{*}\sim\mathcal{U}[-15^{\circ},15^{\circ}]$ and $\psi^{*}\sim\mathcal{U}[-10^{\circ},10^{\circ}]$.
-The robotic agent endeavors to reach the target from the initial state throughout the episode.
-For algorithms incorporating dynamics randomization, a new set of muscle dynamics parameters $\mu$ is sampled at the beginning of each episode by drawing values for each parameter from their respective range.
-Additionally, simulated sensor noise is introduced to the observed state at every action step.
-In the absence of dynamics randomization, muscle parameters are kept constant at their mean values, and no simulated noise is added during the training process.
-To evaluate the effectiveness of the proposed bootstrap and augmentation techniques, we conducted an ablation study on two robotic systems.
-Training results of our learning control design, with and without bootstrap and augmentation, are presented in Fig.
-6.
-The policies were trained and evaluated in simulation with random dynamics.
+The learning rate of DRL algorithms is set to $3\times 10^{-4}$ and the hidden layer size is fixed at $256$.
+The training hyperparameters $N$ and $M$ are empirically chosen based on the system’s complexity.
+Specifically, $N=2000$ and $M=250$ are used for the robotic eye, while $N=3500$ and $M=500$ are used for the robotic wrist.
+When the bootstrap method is not applied, random exploration is employed during the initial $N$ warm-up episodes.
+These hyperparameters are kept consistent across all comparison and ablation studies.
+VI-B Comparison Results Against SOTA Methods
+Several SOTA sample-efficient RL algorithms, designed for state-based observations and continuous action spaces, are selected for comparison.
+REDQ[34] integrates an ensemble of Q-functions with in-target minimization across a random subset of these functions.
+This allows for training with a high Update-To-Data (UTD) ratio, significantly improving sample efficiency.
+DroQ[35] equips the Q-functions with dropout connections and layer normalization techniques, maintaining a high UTD ratio.
+RLPD[36] leverages offline data during online learning by loading demonstrations into the replay buffer and oversampling them during training.
+X-QL[37] directly models soft optimal values and introduces a novel optimization framework based on the Gumbel distribution.
+We trained the proposed SAC-BAR on two robotic systems and compared its performance with these SOTA approaches.
+Performance was measured by the online per-step average reward achieved by the agents, with each algorithm trained across four independent runs.
+Fig.
+6(a) presents the comparative results between SAC-BAR and the selected SOTA RL methods.
+In the robotic eye system, SAC-BAR achieves performance comparable to (and slightly better than) DroQ and RLPD, while significantly outperforming REDQ and X-QL in terms of sample efficiency.
+In the robotic wrist system, SAC-BAR not only significantly outperforms RLPD and DroQ but also demonstrates moderate improvements in convergence speed over REDQ and X-QL.
+Notably, REDQ’s performance relies on a high UTD ratio and an ensemble of Q-functions, resulting in high computational cost; for instance, REDQ requires 18.2 hours for 3,500 training episodes, whereas SAC-BAR completes the same training in only 4.1 hours under our experimental setup.
+Moreover, X-QL exhibited significant instability during the early stages of training in both robotic systems.
+Overall, these results demonstrate that SAC-BAR is better suited than the existing SOTA RL methods for the string-type artificial muscle robots targeted in this study.
+VI-C Ablation Study in Simulation
+To evaluate the effectiveness of the proposed bootstrap and augmentation techniques, an ablation study was conducted on two robotic systems.
+Three variants of our proposed framework were tested as follows: (1) w/o A: This variant excludes the state vector augmentation method.
+(2) w/o B: This variant removes the bootstrap via PID method.
+(3) w/o AB: This baseline variant excludes both mechanisms.
+Each algorithm was trained and evaluated in simulation under randomized dynamics across four independent runs.
+Training results of our proposed SAC-BAR, with and without bootstrap and augmentation, are presented in Fig.
+6(b).
 As illustrated in the figures, our proposed algorithm exhibits the highest level of data efficiency in both robotic systems.
-Compared to the SAC algorithm trained solely with dynamics randomization (referred to as ”No BA”), the integration of bootstrap and augmentation enhances data efficiency and expedites the training process.
-In the robotic eye system, the baseline algorithm achieves an average reward of $-0.10$ after $2000$ training episodes, whereas our approach achieves comparable performance in only approximately $750$ episodes, resulting in a data efficiency increase to $2.67$ times the original level.
-Similarly, in the robotic wrist system, the baseline algorithm achieves an average reward of $-1.61$ after $3500$ episodes, while ours achieves comparable performance in only about $1200$ episodes, leading to a data efficiency increase to $2.92$ times.
-Notably, in the case of the more complex robotic wrist system, the demonstrations provided by the PID controller yield a more significant acceleration effect during the early stages of training.
-Meanwhile, the augmentation method demonstrates a favorable impact on enhancing the training speed in both systems.
-(a) vanilla-SAC
-(b) SAC-R
-(c) SAC-BAR
-(d) Box-and-Whisker Plot
+Compared to the SAC algorithm trained solely with dynamics randomization (referred to as “w/o AB”), the integration of bootstrap and augmentation enhances data efficiency and expedites the training process.
+In the robotic eye system, the baseline algorithm achieves an average reward of $-0.10$ after 2,000 training episodes, whereas our approach achieves comparable performance in only approximately 750 episodes, resulting in a 2.67-fold increase in data efficiency compared to the original level.
+Similarly, in the robotic wrist system, the baseline algorithm achieves an average reward of $-1.61$ after 3,500 episodes, while ours achieves comparable performance in only about 1,200 episodes, leading to a 2.92-fold increase in data efficiency.
+Notably, in the case of the more complex robotic wrist system, we find the demonstrations provided by the PID controller yield a more significant acceleration effect during the early stages of training.
+Meanwhile, the augmentation method demonstrates a favorable impact on enhancing the training speed in both systems and throughout the entire learning process of the algorithms.
+VII Real-World Experiments
+This section presents the real-world experimental results of the proposed SAC-BAR on two testing robotic platforms.
+We first demonstrate the results of an ablation study conducted on the robotic wrist platform, which validate the effectiveness of the proposed dynamics randomization in enhancing control performance in real-world scenarios.
+Next, experimental results on the robotic eye are provided to show the generalizability of SAC-BAR across similar string-type artificial muscle-driven robotic systems.
+Additionally, we investigate the impact of the randomization range on control performance.
+VII-A Ablation Study on Robotic Wrist
+Figure 7: Comparison of experimental trajectories between the baseline and SAC-BAR.
+Target angles are set at ($\theta^{*}=2.5^{\circ},\phi^{*}=-7.5^{\circ}$).
+(a), (b) The system output angles ($\theta,\phi$) and corresponding actions $a_{t}$ of the baseline SAC algorithm.
+(c), (d) Those of our proposed SAC-BAR algorithm.
+To evaluate the impact of each component on real-world control performances, three different control policies are tested on the robotic wrist platform, including the vanilla SAC algorithm, SAC with only randomization of muscles (SAC-R), and our proposed SAC-BAR.
+The parallel robotic wrist starts from the initial state with $\theta=\phi=0^{\circ}$ and attempts to reach and maintain the target angle set $(\theta^{*},\phi^{*})$.
+The action vector $a$ is computed by the trained actors at a time step of $t_{a}=0.5$ s, and each episode spans a duration of $t_{\text{total}}=20$ s.
+In Fig.
+7(a), it is evident that without randomization of muscle dynamics, the vanilla-SAC struggles to achieve the desired orientation.
+While the agent may eventually reach the desired orientation at certain target angles, maintaining that target angle becomes a challenge, resulting in a considerable steady-state error.
+Conversely, as observed in Fig.
+7(c), the proposed SAC-BAR agent achieves the target state in less than $10$ s and consistently remains within its neighborhood thereafter.
+As illustrated by the input action trajectories in Fig.
+7(d), this success is attributed to the agent learning to maximize the action value during the rising phase and promptly reduce it upon approaching the target angles.
+Conversely, in Fig.
+7(b), there exist serious oscillations in the action trajectories of vanilla-SAC during the initial $10$ s, leading to a prolonged rise time and sometimes even failure to reach the target.
 Figure 8: Experimental results of the steady-state error tests in the parallel robotic wrist under selected control algorithms.
-(a), (b), (c) Heat maps illustrating the experimental results for the vanilla SAC algorithm, SAC with the randomization of muscles only (SAC-R), and our complete efficient learning control algorithm (SAC-BAR), respectively.
+(a), (b), (c) Heat maps illustrating the experimental results for the SAC, SAC-R, and SAC-BAR, respectively.
 The combination of target angles $(\theta^{*},\phi^{*})$ is systematically spaced at $2.5^{\circ}$ intervals within the range $(|\theta^{*}|,|\phi^{*}|)\leq 10^{\circ}$.
 The color bars in the visualization indicate the Euclidean distance norm of steady-state errors.
-(d) The box charts illustrating the medians, maxima, minima, quartiles, and outliers of the steady-state errors for the selected comparative control algorithms.
-V-B Ablation Study in Robotic Wrist Experiments
-To evaluate the impact of each component on real-world control performances, three different control policies are tested on the robotic wrist platform, including the vanilla SAC algorithm, SAC with only randomization of muscles (SAC-R), and our proposed efficient learning control framework (SAC-BAR).
-All three algorithms undergo training for an equivalent number of episodes in simulation, employing identical hyperparameters as specified in Table IV.
-During each experimental episode, the parallel robotic wrist starts from the initial state with $\theta=\phi=0^{\circ}$ and attempts to reach and maintain the target angle set $(\theta^{*},\phi^{*})$.
-The action vector $a$ is computed by the trained actors at a time step of $t_{a}=0.5$s, and each episode spans a duration of $t_{\text{total}}=20$s.
-Observations are measured by an IMU positioned at the center of the upper plate.
-TABLE V: Averaged steady-state error over a total of 81 targets in ablation study with the robotic wrist
-Agent Vanilla-SAC SAC-R SAC-BAR
-Mean error (deg) $0.911\pm 0.457$ $0.657\pm 0.386$ $0.580\pm 0.368$
-The experimental trajectories for two pairs of desired target angles, $(\theta^{*},\phi^{*})=(-5^{\circ},7.5^{\circ})$ and $(\theta^{*},\phi^{*})=(2.5^{\circ},-7.5^{\circ})$, of the vanilla SAC algorithm and our proposed SAC-BAR are illustrated in Fig.
-7 and Fig.
-8, respectively.
-In Fig.
-7LABEL:sub@subfig:w_tra1_a and LABEL:sub@subfig:w_tra1_c, it is evident that without randomization of muscle dynamics, the vanilla-SAC agent struggles to achieve the desired orientation.
-While the agent may eventually reach the desired orientation at certain target angles, maintaining that target angle becomes a challenge, resulting in a considerable steady-state error.
-Conversely, as observed in Fig.
-8LABEL:sub@subfig:w_tra2_a and LABEL:sub@subfig:w_tra2_c, the proposed SAC-BAR agent achieves the target state in less than $10$s and consistently remains within its neighborhood thereafter.
-This success is attributed to the agent learning to maximize the action value during the rising phase and promptly reduce it upon approaching the target angles, as illustrated by the input action trajectories in Fig.
-8LABEL:sub@subfig:w_tra2_b and LABEL:sub@subfig:w_tra2_d.
-Additionally, when overshoot occurs, the agent still learns to return to and maintain the target orientation.
-Conversely, in Fig.
-7LABEL:sub@subfig:w_tra1_b and LABEL:sub@subfig:w_tra1_d, there exist serious oscillations in the action trajectories of vanilla SAC agent during the initial $10$s, leading to a prolonged rise time and sometimes even failure to reach the target.
-Figure 10: Experimental trajectories of the foveation test of the 2-DOF robotic eye.
+(d) The box charts illustrating the medians, maxima, minima, and quartiles of the steady-state errors for the selected comparative control algorithms.
+In order to comprehensively evaluate the control performance of the algorithms across different target angles, we designed a steady-state error field test for the robotic wrist.
+We selected a series of target angles $(\theta^{*},\phi^{*})$ spaced at $2.5^{\circ}$ intervals within the defined range $\theta^{*},\phi^{*}\in[-10^{\circ},10^{\circ}]$.
+Each agent was set to run for $25$ s, sequentially targeting each angle pair.
+The steady-state error is measured and averaged over the last $5$s window, determined by computing the Euclidean distance norm between the current and target angles, defined as $e_{\text{ss}}=\sqrt{(\theta-\theta^{*})^{2}+(\phi-\phi^{*})^{2}}$.
+The steady-state tracking errors for different targets of vanilla-SAC, SAC-R, and SAC-BAR are presented in Fig.
+8(a), (b), and (c), respectively.
+The $\mathrm{x}$ and $\mathrm{y}$ axes therein represent $\theta^{*}$ and $\phi^{*}$, while the color bars represent the averaged steady-state errors.
+Following that, box charts are illustrated in Fig.
+8(d).
+These figures demonstrate that vanilla-SAC agent underperforms across nearly the entire defined range, whereas both SAC-R and SAC-BAR show significantly improved control performance.
+The steady-state errors averaged over the defined range, denoted as $\bar{e}_{\text{ss}}$, for three variants and the PID controller are presented in Table III.
+In real-world experiments, all three DRL-based controllers demonstrate superior performance compared to the traditional PID controller, reducing the tracking error by $11.6\%$, $35.9\%$, and $43.7\%$, respectively.
+Compared to the baseline SAC, SAC-R and SAC-BAR achieved reductions in the averaged tracking errors of $27.9\%$ and $36.3\%$, respectively.
+Specifically, our proposed SAC-BAR reaches the desired orientation at every target angle pair with an averaged error of $0.58^{\circ}$.
+Furthermore, as shown in Table III, the observed improvement in real-world performance can be attributed to overcoming the sim-to-real gap, as evidenced by the reduced disparity between simulated and actual results.
+TABLE III: The experimental (Real) and simulation (Sim) results of steady state error using different control methods
+$\bar{e}_{\text{ss}}$ (deg) Control Method
+PID SAC SAC-R SAC-BAR
+Real $1.03\pm 0.49$ $0.91\pm 0.46$ $0.66\pm 0.39$ $0.58\pm 0.37$
+Sim $1.12\pm 0.47$ $0.46\pm 0.27$ $0.47\pm 0.33$ $0.40\pm 0.17$
+The above experimental results indicate that the proposed randomization of muscle dynamics significantly enhances the real-world control performance in the robotic systems of interest, while bootstrap and augmentation techniques contribute relatively less to the improvement.
+We believe that the enhancement observed in SAC-BAR compared to SAC-R primarily stems from the improvement in training efficiency.
+With the incorporation of bootstrap and augmentation methods, SAC-BAR achieves more comprehensive training and higher averaged rewards in simulation after 3,500 training episodes.
+We believe this improvement is potentially more significant in scenarios where exploration is more challenging.
+VII-B Learning Control of Robotic Eye
+Figure 9: Experimental trajectories of the foveation test of the robotic eye.
 Target angles are set at $(\theta^{*}=-5^{\circ},\psi^{*}=7.5^{\circ})$ and $(\theta^{*}=10^{\circ},\psi^{*}=-5^{\circ})$, respectively.
@@
 (b), (d) The corresponding actions $a_{t}$ generated by the trained actor.
-In order to comprehensively evaluate the control performance of the algorithms across different target angles, we design a steady-state error field test for the robotic wrist.
-We select a series of target angles $(\theta^{*},\phi^{*})$ spaced at $2.5^{\circ}$ intervals within the defined range $\theta^{*},\phi^{*}\in[-10^{\circ},10^{\circ}]$.
-Each agent is set to run for $25$s, sequentially targeting each angle pair.
-The steady-state error is measured and averaged over the last $5$s window, determined by computing the Euclidean distance norm between the current and target angles using $e_{\text{ss}}=\sqrt{(\theta-\theta^{*})^{2}+(\phi-\phi^{*})^{2}}$.
-The steady-state errors of the field test of vanilla-SAC, SAC-R, and SAC-BAR are presented in Fig.
-9LABEL:sub@subfig:a, LABEL:sub@subfig:b, and LABEL:sub@subfig:c, respectively.
-The $x$ and $y$ axes therein represent $\theta^{*}$ and $\phi^{*}$, while the color bars represent the averaged steady-state errors.
-Following that, box charts in Fig.
-9LABEL:sub@subfig:d illustrate the medians, maxima, minima, quartiles, and outliers of the steady-state errors for different control policies.
-The figures demonstrate that vanilla-SAC agent performs poorly across nearly the entire defined range, whereas both SAC-R and SAC-BAR outperform significantly.
-The steady-state errors averaged over the defined range for three algorithms are presented in Table V.
-Compared to the baseline vanilla-SAC, SAC-R and SAC-BAR achieved reductions in averaged steady-state errors of $27.9\%$ and $36.3\%$, respectively.
-Specifically, our proposed SAC-BAR reaches the desired orientation at every target angle pair with steady-state errors consistently bounded below $1.5^{\circ}$ except for three outliers.
-The averaged steady-state error of SAC-BAR is $0.58^{\circ}$, demonstrating its superior control performance in accuracy and robustness across different target angles.
-Figure 11: Experimental results in the steady-state error tests of the robotic eye at eyery combination of $(\theta^{*},\psi^{*})$ target angles spaced at $2.5^{\circ}$ intervals for $(|\theta^{*}|,|\psi^{*}|)\leq 10^{\circ}$ where the color bar represents the Euclidean distance norm of steady-state error.
-The above experimental results indicate that randomization of muscle dynamics significantly enhances the real-world control performance of the learning-based controllers in the robotic system of interest, while bootstrap and augmentation techniques contribute less to the improvement.
-We believe that the enhancement observed in SAC-BAR compared to SAC-R primarily stems from the improvement in training efficiency.
-With the incorporation of bootstrap and augmentation methods, SAC-BAR achieves more comprehensive training and higher averaged rewards in simulation after $3500$ training episodes.
-This improvement is potentially more significant in scenarios where training data is limited or exploration proves more challenging.
-V-C Learning Control of Robotic Eye
+Figure 10: Experimental results on the steady-state tracking errors in robotic eye at eyery combination of $(\theta^{*},\psi^{*})$ target angles spaced at $2.5^{\circ}$ intervals for $(|\theta^{*}|,|\psi^{*}|)\leq 10^{\circ}$ where the color bar represents the Euclidean distance norm of steady-state error.
 To demonstrate the applicability of the proposed learning control framework across diverse robotic systems driven by artificial muscle strings, this subsection presents experimental results on an additional platform, the 2-DOF robotic eye.
-We focus on the foveation control in the experiment.
-The robotic eye system, initialized at $\theta=\psi=0^{\circ}$ in ambient room temperature, is configured to foveate at various target angles $(\theta^{*},\psi^{*})$.
+We focus on the foveation control task in the experiment.
+The robotic eye, initialized at $\theta=\psi=0^{\circ}$ at ambient room temperature, is configured to foveate at various target angles $(\theta^{*},\psi^{*})$.
 Considering the system’s rise time, each testing episode lasts for $t_{\text{total}}=15$ s, with an action time step of $t_{a}=0.5$ s.
@@
 These trajectories demonstrate that the measured angles rise to the target in approximately $5$ s, with negligible steady-state error.
-A visual field test is conducted with a series of target angles spaced at $2.5^{\circ}$ intervals within the defined range $|\theta^{*}|,|\psi^{*}|\leq 10^{\circ}$ to evaluate the overall control performance of the algorithm.
+A similar field test is conducted with a series of target angles spaced at $2.5^{\circ}$ intervals within the defined range $|\theta^{*}|,|\psi^{*}|\leq 10^{\circ}$ to evaluate the overall control performance of the algorithm.
 The robotic eye is tested to foveate at each target in sequence.
 The results, illustrated in Fig.
-11, reveal excellent control performance, with steady-state errors consistently below $1^{\circ}$ for most target angles.
-In prior research [23], we adopted a reinforcement learning-based controller using deep deterministic policy gradient (DDPG) algorithm [50], trained in simulation for $50000$ episodes.
-Additionally, a reduced-order state observer estimated the real-time temperature of SCP muscles, resulting in satisfactory control performance.
-In this paper, our proposed algorithm demonstrates comparable control performance with just $2000$ training episodes, representing only $1/25$ of the original requirement.
-Remarkably, it accomplishes this without the need for real-time muscle temperature information.
-VI Conclusions
-This paper presented a learning-based control framework designed for the string-type artificial muscle-driven robotic systems, leveraging the SAC algorithm with a GRU hidden layer as the baseline DRL algorithm.
-To enhance training data efficiency and expedite convergence speed, we proposed two methods including bootstrap by PID and state-based data augmentation.
+10, reveal solid control performance, with steady-state errors consistently below $1^{\circ}$ for most target angles and an average of $0.56^{\circ}$.
+In our prior research[20], we employed a DRL-based controller using the Deep Deterministic Policy Gradient algorithm[43], which was trained in simulation for 50,000 episodes.
+Additionally, a reduced-order state observer was used to estimate the real-time temperature of SCP muscles, achieving satisfactory control performance.
+In this paper, our proposed algorithm achieves comparable control performance with only 2,000 training episodes, representing just 1/25 of the original training time.
+Moreover, this is accomplished without the need to estimate real-time muscle temperature information under the condition of ambient temperature remaining relatively stable.
+VII-C Randomization Ranges on Control Performance
+TABLE IV: Comparison of control errors between models trained under different randomization ranges
+$\bar{e}_{\text{ss}}$ (deg) Randomization Range
+Small Original Large
+Real $0.63\pm 0.49$ $0.58\pm 0.37$ $0.76\pm 0.43$
+Sim $0.38\pm 0.17$ $0.40\pm 0.17$ $0.74\pm 0.41$
+In designing dynamics randomization, maintaining a balance between randomization level and model fidelity is crucial.
+Excessive randomization leads to an overly conservative controller, which may over-generalize and fail to leverage the more predictable aspects of system dynamics.
+Conversely, insufficient randomization causes the controller to overfit to a narrow set of conditions, thus limiting its performance when deployed in real-world systems.
+To investigate this trade-off, we conducted an experiment in which our control framework was trained with two additional randomization ranges—“small” and “large.” The small range was defined as half of the original randomization range, while the large range was doubled.
+We evaluated the performance of controllers trained under these varying randomization ranges in both simulated and real-world environments on the robotic wrist platform.
+The results, as summarized in Table IV, indicate that the control performance experiences a significant decline when randomization range is excessively large.
+Conversely, the smaller randomization range impairs the controller’s generalization capability from simulation to reality.
+Overall, a balanced range proves essential for ensuring robust performance.
+VIII Conclusion
+This paper presents a learning-based control framework SAC-BAR designed for the string-type artificial muscle-driven robotic systems.
+To enhance training data efficiency and expedite convergence speed, we proposed two methods, namely bootstrap via PID and state-based data augmentation.
 Additionally, we incorporated randomization of artificial muscle dynamic parameters during training to address the sim-to-real gap for enhancing real-world control performance.
-Extensive experiments and ablation studies on two systems—a 2-DOF robotic eye driven by SCP actuators and a parallel robotic wrist driven by TCA muscles—validated the effectiveness and generalizability of the proposed control framework.
-In simulation, the designed bootstrap and augmentation methods significantly reduced the training episodes required for DRL-based controllers.
-Real-world application experiments demonstrated that randomization of muscle dynamics notably enhanced the control performance of learning-based controllers.
-Furthermore, our framework exhibited satisfactory performance on both robotic platforms, indicating its great potential of applications in similar systems driven by artificial muscle strings.
-Despite the notable performance of the proposed algorithm, several limitations exist.
-Noticeable jitters in action trajectories may prolong system rise time.
-Additionally, the configuration of the reward function still has a considerable impact on the experimental outcomes.
-In future work, we plan to analyze these issues comprehensively using extensive experimental data and refine our algorithms to effectively address these challenges.
+Extensive experiments and ablation studies on two systems—a robotic eye driven and a parallel robotic wrist—validated the effectiveness and generalizability of the proposed control framework.
+In simulation, the designed bootstrap and augmentation methods reduced the required training episodes by approximately 60% compared to the baseline algorithm.
+Real-world experiments demonstrated that randomization of muscle dynamics improved the control performance of DRL-based controllers in targeted robotic systems, reducing the average error by 36.3%.
+Furthermore, our framework showed satisfactory performance on both robotic platforms, highlighting its potential applicability to similar artificial muscle robots.
+Despite the strong performance of the proposed SAC-BAR, several limitations exist.
+The configuration of the reward function has a significant impact on experimental outcomes, often requiring empirical tuning to ensure optimal controller performance.
+The dynamic models of the artificial muscles and robotic wrist are simplified and do not fully represent their true dynamics.
+In future work, we plan to explore automated methods for reward function tuning, potentially incorporating adaptive or self-tuning mechanisms that respond dynamically to performance metrics.
+To better account for real-world variability, we will extend the randomization of muscle dynamics by incorporating high-fidelity models that capture nonlinearity and time-dependent factors, using data-driven or hybrid modeling approaches.

### v4

@@
 Index Terms:
-Reinforcement learning for robot control, artificial muscle, sim-to-real, sample efficiency.
+Reinforcement learning for robot control, artificial muscle, sample efficiency, sim-to-real.
 I Introduction
@@
 Well known for their high power-to-weight ratios and inherent compliance[1], artificial muscles offer a unique combination of strength and flexibility conducive to versatile and adaptive robot designs.
-Over the past decade, robots driven by artificial muscles have achieved significant advancements, demonstrating impressive performance across diverse domains, including but not limited to biomimetic robots[2, 3], robot manipulators[4, 5], soft robots[6, 7], medical robot[8], origami robots[9], and marine robots[10].
-Previous research efforts have primarily centered around structural design and open-loop motion performance evaluation[2, 3, 4, 5, 6, 7, 8, 9, 10].
+Over the past decade, robots driven by artificial muscles have achieved significant advancements, demonstrating impressive performance across diverse domains, including but not limited to biomimetic robots[2, 3], robot manipulators[4, 5], and soft robots[6, 7].
+Previous research efforts have primarily centered around structural design and open-loop motion performance evaluation[2, 3, 4, 5, 6, 7].
 Despite their potential, achieving accurate control of string-type artificial muscle-driven robots remains challenging due to their highly nonlinear and complex dynamics.
@@
 The DRL algorithms are then trained within this simulation environment to develop an effective control policy.
-Finally, the learned control policy is transferred to the physical robotic platform to achieve the desired control objectives.
+Finally, the learned policy is transferred to the physical robotic platform to achieve desired control performance.
 Figure 1: Schematic of standard procedure of applying DRL control in artificial muscle-driven robotic systems.
@@
 Another challenge is the low learning efficiency of DRL algorithms, which require substantial amounts of training data and iterations due to the complexity of neural networks and the trial and error learning process[21].
-Addressing these two challenges is crucial for advancing control technologies in artificial muscle-driven robots, which are increasingly employed in soft robotics and bio-inspired systems.
+Addressing these two challenges is crucial for advancing control technologies in artificial muscle-driven robots.
 This paper presents a DRL-based control strategy designed to enhance the data efficiency and real-world applicability of controllers for artificial muscle-driven robots.
@@
 Additionally, we leverage a zero-shot sim-to-real transfer technique on artificial muscle dynamics.
-The experimental results demonstrate a significant improvement in real-world control performance with the averaged error reduced by $36.3\%$ through the application of this method.
-The remainder of this paper is organized as follows.
-Section II provides a review of the relevant literature.
-Section III introduces the basic concepts of applying reinforcement learning to the robotic systems studied.
-Section IV offers a concise overview of the robotic platforms utilized in this study.
-In Section V, we present a detailed description of our proposed efficient learning control framework.
-Section VI and Section VII discuss the experimental results, including training outcomes in simulation and control performances in real-world scenarios, respectively.
-Finally, concluding remarks and current limitations are discussed in Section VIII.
+The experimental results demonstrate a significant improvement in real-world control performance with the averaged error reduced by $36.3\%$.
+The main contributions of this paper are threefold.
+First, we enhance the training of DRL by leveraging PID controller-generated trajectories to accelerate early-stage training and introducing a novel state-based augmentation method that transforms single data transitions into multiple augmented samples.
+Second, we propose a zero-shot sim-to-real transfer technique specifically tailored for artificial muscle dynamics, effectively addressing the sim-to-real gap caused by dynamic model inaccuracies and parameter variability.
+Third, we evaluate our approach through extensive experiments on two distinct robotic systems, demonstrating significant reductions in both training data requirements and real-world control errors.
+These results validate the effectiveness of the proposed method and highlight its potential applicability to similar systems.
 II Related Work
@@
 Many of these studies have noted that the low sample efficiency often leads to prolonged training time[17, 18], but few have focused specifically on improving training efficiency.
-Several existing state-of-the-art (SOTA) reinforcement learning (RL) methods may be applicable[34, 35, 36, 37], but they have not been evaluated within the context of the robotic systems we are addressing.
+Several existing state-of-the-art (SOTA) data efficient reinforcement learning (RL) methods may be applicable[31, 32, 33, 34], but they have not been evaluated within the context of the robotic systems we are addressing.
 The unique nonlinear and slow response characteristics of artificial muscle may present challenges that are not adequately addressed by these existing methods.
@@
 Continuous rapid transitions between heating and cooling often cause irreversible damage to the muscles, making direct training on real robots unsafe.
-It is worth noting that in recent years, there have been some efforts to bridge the sim-to-real gap of artificial muscles.
-Many researchers[22, 23, 39] have proposed more accurate yet significantly more complex dynamic models.
+Many researchers[19, 20, 36] have proposed more accurate yet significantly more complex dynamic models to bridge the sim-to-real gap of artificial muscles.
 In this paper, however, we aim to directly improve the performance of DRL-based controllers by employing a sim-to-real transfer technique, thereby reducing the need for complicated mathematical modeling and precise parameter identification for artificial muscles.
@@
   • $\mathcal{R}$ is the reward function that maps the current state $s_{t}$ and action $a_{t}$ to an immediate reward $r_{t}=\mathcal{R}(s_{t},a_{t})$.
-The design of reward function plays a vital role in the learning process of the DRL agents, which learn the optimal policy under the guidance thereby.
 In this paper, we develop a generalized quadratic-form reward function designed for the robotic systems of interest:
     $r_{t}=-(e^{\top}_{t}Q_{e}e_{t}+a^{\top}_{t}R_{a}a_{t})+r_{\text{bonus}},$ (1)
-    where $e_{t}=|\bm{y}^{*}-\bm{y}_{t}|$ is the system tracking error, and $Q_{e}$ and $R_{a}$ represent the output and control weight matrices, respectively.
-A bonus reward $r_{\text{bonus}}$ is given when the system tracking error falls below a sufficiently small threshold.
+    where $e_{t}=|\bm{y}^{*}-\bm{y}|$ is the system tracking error, and $Q_{e}$ and $R_{a}$ represent the output and control weight matrices, respectively.
+These weights are determined based on empirical considerations with diagonal elements set to ensure terms in the reward function have comparable magnitudes.
+A bonus reward $r_{\text{bonus}}$ is given when the tracking error falls below a sufficiently small threshold.
   • $f$ represents the state transition function and $\gamma$ is the discount factor, which is set to $0.99$ in our experiments.
@@
 This section presents an overview of the actuators and robotic platforms used in this paper, including SCP and TCA artificial muscles (Fig.
-2), a 2-DOF robotic eye (Fig.
+2), a two-degrees-of-freedom (DOF) robotic eye (Fig.
 3) and a parallel robotic wrist (Fig.
 4).
-For each system, we briefly describe the hardware design, state space, action space, and reward function.
+For each system, we briefly describe the hardware design, state space, action space, and reward function design.
 IV-A Artificial Muscles
@@
 3, the hemispherical eyeball is driven by two pairs of SCP muscle actuators, accommodating a Raspberry Pi Mini Camera (RPiMC).
-The primary control objective is to precisely and rapidly adjust the eyeball’s orientation towards the desired target angles by regulating the SCP muscles, namely foveation control in the eye movement study.
+Here, $Q_{i}$ represents the anchor position of muscle $m_{i}$ on the eyeball, through which the force generated by SCP muscle is transmitted.
+$P_{i}$ denotes the fixed endpoint of muscle $m_{i}$.
+The primary control objective is to precisely and rapidly adjust the eyeball’s orientation towards the desired target angles, namely foveation control in eye movement study.
 The eyeball’s center of rotation, denoted as $\bm{O}\in\mathbb{R}^{3}$, remains fixed at the origin within the reference frame $O_{\mathrm{xyz}}$.
@@
 where $\theta,\dot{\theta},\psi,\dot{\psi},$ represent the pitch, yaw angles, and their respective angular velocities for control.
-$\Delta T_{i}$ represents the temperature difference between the ambient temperature $T_{0}$ and the current temperature $T_{i}$ of muscle $m_{i}$.
+$\Delta T_{i}$ represents the temperature difference between the ambient temperature $T_{\text{amb}}$ and the current temperature $T_{i}$ of muscle $m_{i}$.
 Figure 4: The parallel robotic wrist prototype actuated by three TCA muscles.
@@
 The Euler angles and corresponding angular velocities are measured through visual feedback from the embedded RPiMC.
-The actuating voltage applied to muscle $m_{i}$ is denoted as $V_{i}$.
+In this paper, the actuating voltage applied to muscle $m_{i}$, denoted as $V_{i}$, is treated as a continuous variable within the defined range.
 Given the antagonistic configuration of the muscle pairs, the control input is defined by a constrained action vector, i.e.,
@@
 The constraint requires that only one muscle is activated at each time step in the pairs $(m_{1},m_{2})$ and $(m_{3},m_{4})$.
-The reward function follows the structure of (1), where the output and control weight matrices were set to $Q_{e}=\text{diag}\{0.05,0.25,0.05,0.25\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively.
-They were chosen based on empirical tuning to balance the trade-off between output error and control effort.
-The bonus reward is granted when the tracking error is sufficiently small, given by
+The reward function follows the structure of (1), where the output and control weight matrices are set to $Q_{e}=\text{diag}\{0.05,0.25,0.05,0.25\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively.
+The bonus reward is given by
 $r_{\text{bonus}}=2\times\left(\mathds{1}\{|\psi-\psi^{*}|<0.3\}+\mathds{1}\{|\theta-\theta^{*}|<0.3\}\right).$ (5)
@@
 Research into artificial muscle-driven robotic wrist joints is of great importance to several domains, including the development of dexterous robotic systems with human-like capabilities, advancements in rehabilitation and prosthetics technology, and the enhancement of robotic manipulation versatility across various fields.
-We designed a parallel robotic wrist prototype driven by three equally spaced compound TCAs[27] as illustrated in Fig.
+We design a parallel robotic wrist prototype driven by three equally spaced compound TCAs[24] as illustrated in Fig.
 4.
@@
 Linkages between the two plates are constructed to support a two degree-of-freedom motion of the upper plate[37].
-Given this configuration, we developed a simplified dynamics model under the assumption of a fixed rotational center for the upper plate.
+Given this configuration, we develop a simplified dynamics model under the assumption of a fixed rotational center for the upper plate.
 Similar to the foveation control of the robotic eye, the control objective entails precise regulation of the moving upper plate’s orientation.
-This necessitates adjusting the applied actuating voltages across the three muscle strings to drive the upper plate from its initial equilibrium state to predefined target angles.
 The reference frame $O_{\mathrm{xyz}}$ is fixed on the moving upper plate, with the $\mathrm{z}$-axis perpendicular to the upper plate and pointing outward.
@@
 where $\theta^{*}$ and $\phi^{*}$ denote the target angles, with the respective Euler angles and angular velocities measured using the inertial measurement unit (IMU) positioned at the center of the upper plate.
-Additionally, the input action vector $a_{t}$ of the system is defined as:
+The input action vector $a_{t}$ of the system is defined as:
 […]
-The output and control weight matrices of the reward function are set to $Q_{e}=\text{diag}\{0.05,0.2,0.05,0.2\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively.
-When the tracking error is sufficiently small, a bonus reward is granted, i.e.,
+The weight matrices are set to $Q_{e}=\text{diag}\{0.05,0.2,0.05,0.2\}$ and $R_{a}=\text{diag}\{0.01,0.01,0.01\}$, respectively. The bonus reward is given by
 $r_{\text{bonus}}=2\times\left(\mathds{1}\{|\theta-\theta^{*}|<0.5\}+\mathds{1}\{|\phi-\phi^{*}|<0.5\}\right).$ (9)
+The threshold values in (5) and (9) are selected empirically based on testing experiments.
+Lower values may result in slower convergence and lower training stability due to reduced positive reinforcement, while higher values could diminish the precision of the learned control policy.
 V Learning Control Framework
@@
 Subsequently, the agent interacts directly with the simulation environment to update its policy.
-All transition data collected from the environment undergoes a state-based augmentation strategy before being stored into the replay buffer.
-To mitigate the sim-to-real gap, we introduce a randomization method on the artificial muscle dynamics at the beginning of each training episode.
-Eventually, the trained actors (control policies) can be directly implemented in real-world robotic systems to achieve the desired control performance.
+All transition data collected from the environment undergoes a state-based augmentation strategy before being stored into replay buffer.
+To mitigate the sim-to-real gap, we introduce a randomization method on the artificial muscle dynamics.
+In this paper, we denote the bootstrap method by letter ‘B’, the augmentation method by letter ‘A’, and the randomization of muscle dynamics by letter ‘R’.
+Thus, our proposed SAC algorithm with all three improvements is denoted by SAC-BAR.
 V-A Soft Actor-Critic
-Soft Actor-Critic (SAC) [29], a state-of-the-art DRL algorithm, is widely adopted as a strong baseline for continuous control tasks in robotics due to its robustness, sample efficiency, and model-free nature.
-Unlike traditional DRL algorithms, SAC incorporates an entropy-maximizing approach, with its learning objective defined as:
+SAC [26] is widely adopted as a strong baseline for continuous control tasks in robotics due to its robustness and sample efficiency.
+SAC incorporates an entropy-maximizing approach, with its learning objective defined as:
 $J(\pi)=\sum_{t=0}^{T}\mathbb{E}_{(s_{t},a_{t})\sim\rho_{\pi}}\left[\mathcal{R}(s_{t},a_{t})+\alpha\mathcal{H}(\pi(\cdot|s_{t}))\right],$ (10)
-where $\mathcal{H}(\pi(\cdot|s_{t}))$ represents the entropy of the policy $\pi$, and $\alpha$ serves as a temperature parameter.
+where $\mathbb{E}_{(s_{t},a_{t})\sim\rho_{\pi}}$ represents the expectation over the state-action pairs $(s_{t},a_{t})$ sampled from the policy distribution $\rho_{\pi}$, $\mathcal{H}(\pi(\cdot|s_{t}))$ represents the entropy of the policy $\pi$, and $\alpha$ serves as a temperature parameter.
 This entropy-regularized objective improves the robustness of DRL and accelerates its training [21].
-To process sequential information critical for addressing partial observability issue and implementing dynamics randomization, we integrate a Gated Recurrent Unit (GRU) [41] as the hidden layer in the actor-critic architecture.
+We integrate a Gated Recurrent Unit (GRU) [38] as the hidden layer in the actor-critic architecture to process sequential information.
 V-B Randomization of Muscle Dynamics
@@
 […]
-$c$ $(\text{N}/^{\circ}\text{C})$ $0.0055$ uniform$(0.85,1.15)$
-$C_{\text{th}}$ $(\text{Ws}/^{\circ}\text{C})$ $0.28$ uniform$(0.8,1.2)$
-$\lambda$ $(\text{W}/^{\circ}\text{C})$ $0.094$ uniform$(0.85,1.15)$
-$R$ $(\Omega)$ $20.0$ uniform$(0.9,1.1)$
-$\theta,\psi$ noise $(^{\circ})$ normal$(0,0.1)$
-$\dot{\theta},\dot{\psi}$ noise $(^{\circ}/\text{s})$ normal$(0,0.05)$
+$c$ $(\text{N}/^{\circ}\text{C})$ $0.0055$ $\mathrm{uniform}(0.85,1.15)$
+$C_{\text{th}}$ $(\text{Ws}/^{\circ}\text{C})$ $0.28$ $\mathrm{uniform}(0.8,1.2)$
+$\lambda$ $(\text{W}/^{\circ}\text{C})$ $0.094$ $\mathrm{uniform}(0.85,1.15)$
+$R$ $(\Omega)$ $20.0$ $\mathrm{uniform}(0.9,1.1)$
+$\theta,\psi$ noise $(^{\circ})$ $\mathrm{normal}(0,0.1)$
+$\dot{\theta},\dot{\psi}$ noise $(^{\circ}/\text{s})$ $\mathrm{normal}(0,0.05)$
 The dynamic model of the artificial muscles considered in this paper is based on the literature[23], where the muscle’s elasticity behavior is linearized to establish a dynamical system comprising a mechanical model characterized by stiffness and damping, given by
 $F=k(x-x_{0})+b\dot{x}+c(T(t)-T_{\text{amb}}),$ (11)
-where $x$ and $x_{0}$ denote the loaded length and unloaded length of the actuator, respectively, while $k$, $b$, and $c$ denote the mean stiffness, damping, and temperature effect coefficients of the actuator, respectively.
+where $x$ and $x_{0}$ denote the current length and resting length of the actuator, respectively, while $k$, $b$, and $c$ denote the mean stiffness, damping, and temperature effect coefficients of the actuator, respectively.
 $T(t)$ denotes the current actuator temperature, while $T_{\text{amb}}$ stands for the ambient temperature of the environment.
@@
 where $C_{\text{th}}$ represents the thermal mass of the actuator (measured in $\text{Ws}/^{\circ}\text{C}$), $P(t)$ indicates the heat applied to the actuator, $\lambda$ is the absolute thermal conductivity of the actuator with respect to the ambient environment (in $\text{W}/^{\circ}\text{C}$).
-The applied power, $P(t)$, is controlled through the voltage applied to the muscle, expressed as $P(t)=V(t)^{2}/R$.
+The applied power, $P(t)$, is controlled through the voltage applied to the muscle, expressed as $P(t)=V(t)^{2}/R$, where $R$ is the resistance of the artificial muscle.
+Equations (11) and (12) describe the dynamics of the artificial muscles.
+In the simulation, when the RL agent generates an action $a_{t}$, representing the voltages applied to the muscles, these equations are numerically solved to update the muscle states and compute their output forces.
+The computed forces are then incorporated into the dynamic model of the robotic system to update the overall system state ($s_{t}$).
 TABLE II: Dynamics parameters of randomization and their respective ranges in the robotic wrist
@@
 […]
-$c$ $(\text{N}/^{\circ}\text{C})$ $0.0707$ uniform$(0.85,1.15)$
-$C_{\text{th}}$ $(\text{Ws}/^{\circ}\text{C})$ $3.06$ uniform$(0.8,1.2)$
-$\lambda$ $(\text{W}/^{\circ}\text{C})$ $0.1189$ uniform$(0.85,1.15)$
-$R$ $(\Omega)$ $10.0$ uniform$(0.9,1.1)$
-$\theta,\phi$ noise $(^{\circ})$ normal$(0,0.1)$
-$\dot{\theta},\dot{\phi}$ noise $(^{\circ}/\text{s})$ normal$(0,0.05)$
+$c$ $(\text{N}/^{\circ}\text{C})$ $0.0707$ $\mathrm{uniform}(0.85,1.15)$
+$C_{\text{th}}$ $(\text{Ws}/^{\circ}\text{C})$ $3.06$ $\mathrm{uniform}(0.8,1.2)$
+$\lambda$ $(\text{W}/^{\circ}\text{C})$ $0.1189$ $\mathrm{uniform}(0.85,1.15)$
+$R$ $(\Omega)$ $10.0$ $\mathrm{uniform}(0.9,1.1)$
+$\theta,\phi$ noise $(^{\circ})$ $\mathrm{normal}(0,0.1)$
+$\dot{\theta},\dot{\phi}$ noise $(^{\circ}/\text{s})$ $\mathrm{normal}(0,0.05)$
 The core idea behind dynamics randomization is to introduce randomness into specific dynamic parameters within the simulator during training.
-This approach allows the DRL agent to learn a policy adaptable to various dynamics conditions, including those encountered in real-world experiments.
-Specifically, we randomize several critical muscle parameters within the simulated dynamics environment.
-A distribution $\rho_{\mu}$ of the muscle parameter set $\mu$ is established to encompass the variability of muscle dynamic parameters for dynamics randomization purposes.
-This approach is expected to be particularly effective in our setup, as artificial muscles are known to exhibit significant variability due to factors such as material degradation, slight manufacturing differences, and environmental temperature changes.
-According to the muscle models (11) and (12), the dynamic parameter set $\mu$ for randomization is given by
-$\mu:=\left\{k,b,c,C_{\text{th}},\lambda,R\right\}.$ (13)
-Table I and II outline the distribution $\rho_{\mu}$ of SCP and TCA muscles, respectively, along with the observation noise added for the robotic eye and wrist systems.
-$\rho_{\mu}$ is centered on the mean values derived from system identification to align the data distribution in simulation more accurately with that in reality.
-To strike the right balance between randomization and fidelity, we carefully tune the range and distribution of randomized parameters based on empirical observations of muscle behavior in real-world scenarios.
-Specifically, the range of each parameter is determined based on the distribution observed across different real muscle strings, as outlined in [26].
-This approach allows us to maintain a high degree of fidelity while simultaneously enabling the controller to generalize effectively to real-world uncertainties.
-Details on the impact of varying randomization ranges on experimental results are discussed in Section VII-C.
-During the training process, the physical parameters of the actuators are randomized within predefined scaling ranges (i.e., $\mu\sim\rho_{\mu}$) at the start of each training episode and remain constant throughout.
-Additionally, Gaussian noise is added to the policy’s observation inputs to simulate the noise typically present in real-world sensor data.
-This observation noise varies at each time step and across different state variables.
-Although this study employs a simplified dynamic model for artificial muscles, the proposed method can be extended to more complex muscle dynamics models to further enhance control performance.
+While this method is typically used in the field of manipulation and locomotion[39, 40], its integration into the control of artificial muscle-driven robots is largely underexplored.
+We hypothesize that this method is particularly effective due to the unique uncertainties inherent in these systems.
+Artificial muscle actuators are subject to significant parameter variability caused by factors such as manufacturing tolerances, wear, and environmental conditions.
+By exposing the RL agent to a diverse range of dynamics during training, this approach facilitates the development of robust control policies that generalize effectively to real-world conditions with unmodeled dynamics.
+Specifically, a distribution $\rho_{\mu}$ of the muscle parameter set $\mu$ is established to encompass the variability of muscle dynamic parameters for dynamics randomization purposes.
+According to (11) and (12), the dynamic parameter set $\mu$ for randomization is given by $\mu:=\left\{k,b,c,C_{\text{th}},\lambda,R\right\}$, where :=assign:=:= is used to denote an assignment operation in this paper.
+Table I and II outline the distribution $\rho_{\mu}$ of SCP and TCA muscles, respectively, along with the observation noise added for the robotic eye and wrist systems, where $\mathrm{uniform}(a,b)$ denotes a uniform distribution over interval $[a,b]$ and $\mathrm{normal}(a,b)$ denotes a normal distribution with mean $a$ and standard deviation $b$.
+The distribution $\rho_{\mu}$ is centered on the mean values derived from system identification to ensure alignment with the physical systems.
+The randomization ranges of these parameters are determined empirically to capture realistic variability while ensuring physical feasibility.
+These ranges are derived from standard deviations reported in prior studies on artificial muscle properties, including influences from material properties, environmental factors, and manufacturing differences[23, 25, 19].
+Details on the impact of randomization variance on control performance are discussed in Section VII-C.
+During the training process, the physical parameters of the actuators are randomized within predefined scaling ranges (i.e., $\mu\sim\rho_{\mu}$) at the start of each episode and remain constant throughout.
+Observation noise is sampled independently at each time step from a Gaussian distribution with a constant mean and variance, as specified in Table I and II.
 V-C Bootstrap via PID
@@
 Within the first $M$ episodes of training, the PID controller interacts with the environment and generates trajectories, which are then stored in the replay buffer along with the rewards.
-Unlike traditional behavior cloning methods, we do not employ a supervised loss for demonstration learning, allowing the agent to avoid strictly mimicking the PID control policy.
-Subsequently, the agent continues interacting with the environment to refine its policy, eventually surpassing the PID controller and achieving superior control performance.
+Compared to using explicit pre-training with PID data or a supervised loss for demonstration learning, our design mitigates the potential risk of the policy overfitting to suboptimal PID behavior before transitioning to RL.
 V-D Augmentation of State Vector
-Data augmentation refers to a class of techniques aimed at increasing the size and improving the quality of training datasets, thereby facilitating the development of more robust deep learning models[42].
-The fundamental concept of data augmentation involves transforming limited data into diverse and functionally equivalent datasets.
+Data augmentation refers to a class of techniques aimed at increasing the size and improving the quality of training datasets[41].
+The concept involves transforming limited data into diverse and functionally equivalent datasets.
 Given the labor-intensive and computationally expensive nature of collecting training data in robotic systems, we integrate data augmentation into the agent’s learning process.
@@
 5:  Initialize dynamics parameters distribution $\rho_{\mu}$
-6:  for $\text{step}=1$ to $M$ do
+6:  for $\text{episode}=1$ to $M$ do
 7:     Sample dynamics $\mu\sim\rho_{\mu}$
@@
 14:
-15:  for $\text{step}=M+1$ to $N$ do
+15:  for $\text{episode}=M+1$ to $N$ do
 16:     Sample dynamics $\mu\sim\rho_{\mu}$
@@
 32:  end for
-As commonly acknowledged, rewards serve as pivotal guides in the learning process of DRL agents.
-Therefore, our objective is to generate additional data transitions with different rewards.
-As described in Section III, the reward $r_{t}$ is calculated as:
+Our objective is to generate additional data transitions with different rewards. As described in Section III, the reward $r_{t}$ is calculated as:
 $r_{t}=\mathcal{R}(\bm{x}_{t},\bm{y}^{*},a_{t}).$ (13)
@@
 which creates different transition tuples at a negligible computational cost.
-In this paper, we denote the bootstrap method by the letter ’B’, augmentation method by the letter ’A’, and the randomization of muscle dynamics by the letter ’R’.
-Thus, our proposed SAC algorithm with all three improvements is denoted by SAC-BAR.
-The pseudo code of our overall SAC-BAR learning control framework is presented in Algorithm 1.
+We choose to perturb the target vectors instead of the state vectors because perturbing state vectors may introduce physical inconsistencies, which can prevent the agent from learning accurate dynamics.
+The pseudo code of SAC-BAR learning control framework is presented in Algorithm 1.
 Figure 6: Reward trajectories with respect to the training episode number.
-(a) Comparison results between our SAC-BAR and SOTA methods including REDQ, DroQ, RLPD, and X-QL.
-(b) Results of ablation study where ‘B’, ‘A’, and ‘R’ indicate bootstrap, augmentation, and randomization techniques, respectively.
+The shaded regions indicate the standard deviation of the rewards across multiple runs.
+(a) Comparative results between our proposed SAC-BAR method and SOTA approaches, including REDQ, DroQ, RLPD, and X-QL.
+(b) Comparative results from the ablation study, where ‘A’, ‘Ps’, ‘B’, and ‘Pt’ represent the augmentation method, state perturbation variation, bootstrap mechanism, and pre-training variation, respectively.
 VI Simulation Experiments
-In this section, we present the training results from simulation environments, comparing our proposed SAC-BAR against several SOTA reinforcement learning algorithms in terms of sample efficiency on two robotic systems.
-Additionally, we perform ablation studies to evaluate the effectiveness of the proposed bootstrap and augmentation methods.
+This section presents the training results from simulation environments, comparing our proposed SAC-BAR against several SOTA reinforcement learning algorithms in terms of sample efficiency on two robotic systems.
+Additionally, we perform ablation studies to evaluate the effectiveness of the proposed methods and discuss the choice of hyperparameters.
 VI-A Training Setup
-At the beginning of each training episode, a target state is randomly selected from a uniformly distributed range.
-For instance, in the robotic eye scenario, the target angles adhere to $\theta^{*}\sim\mathcal{U}[-15^{\circ},15^{\circ}]$ and $\psi^{*}\sim\mathcal{U}[-10^{\circ},10^{\circ}]$.
-The robotic agent strives to reach the target from the initial state throughout the episode.
-In this paper, simulation environments are all written in Python.
-Training and testing in simulation are conducted on a computer with an Intel Xeon Platinum 8383C CPU and an NVIDIA GeForce RTX 4090 GPU.
-The algorithms are implemented in PyTorch and trained using the Adam optimizer.
-The learning rate of DRL algorithms is set to $3\times 10^{-4}$ and the hidden layer size is fixed at $256$.
-The training hyperparameters $N$ and $M$ are empirically chosen based on the system’s complexity.
+In each training episode, the robotic agent aims to reach a target state randomly selected from a uniformly distributed range.
+The initial PID controller parameters are set as $K_{p}=2.1$, $K_{i}=0.2$, and $K_{d}=0.5$ for the robotic eye, and $K_{p}=3.3$, $K_{i}=0.5$, and $K_{d}=0.3$ for the robotic wrist.
+The number of data augmentation is set as $n=10$.
+The neural network architecture for both the actor and critic consists of a fully connected input layer, a GRU layer with 256 hidden units, and a fully connected output layer, with ReLU as the activation function.
+The model is trained using a batch size of 20 trajectories and an experience replay buffer with a capacity of $1\times 10^{5}$ trajectories.
+The learning rate is set to $3\times 10^{-4}$.
+Additionally, the SAC temperature parameter $\alpha$ is configured to be auto-tuned, as described in[26].
+The hyperparameters $N$ and $M$ are empirically determined based on the system’s complexity.
 Specifically, $N=2000$ and $M=250$ are used for the robotic eye, while $N=3500$ and $M=500$ are used for the robotic wrist.
 When the bootstrap method is not applied, random exploration is employed during the initial $N$ warm-up episodes.
-These hyperparameters are kept consistent across all comparison and ablation studies.
+The algorithms are implemented in PyTorch and trained on a computer with an Intel Xeon Platinum 8383C CPU and an NVIDIA GeForce RTX 4090 GPU.
 VI-B Comparison Results Against SOTA Methods
-Several SOTA sample-efficient RL algorithms, designed for state-based observations and continuous action spaces, are selected for comparison.
-REDQ[34] integrates an ensemble of Q-functions with in-target minimization across a random subset of these functions.
-This allows for training with a high Update-To-Data (UTD) ratio, significantly improving sample efficiency.
-DroQ[35] equips the Q-functions with dropout connections and layer normalization techniques, maintaining a high UTD ratio.
-RLPD[36] leverages offline data during online learning by loading demonstrations into the replay buffer and oversampling them during training.
-X-QL[37] directly models soft optimal values and introduces a novel optimization framework based on the Gumbel distribution.
-We trained the proposed SAC-BAR on two robotic systems and compared its performance with these SOTA approaches.
-Performance was measured by the online per-step average reward achieved by the agents, with each algorithm trained across four independent runs.
+Several SOTA sample-efficient RL algorithms, tailored for state-based observations and continuous action spaces, are selected for comparison.
+These include REDQ[31], DroQ[32], RLPD[33], and X-QL[34], which represent diverse approaches to improving sample efficiency.
+We train SAC-BAR on two robotic systems and compare its performance with these SOTA approaches.
+Performance is measured by the online per-step average reward achieved by the agents, with each algorithm trained across four independent runs.
 Fig.
@@
 Notably, REDQ’s performance relies on a high UTD ratio and an ensemble of Q-functions, resulting in high computational cost; for instance, REDQ requires 18.2 hours for 3,500 training episodes, whereas SAC-BAR completes the same training in only 4.1 hours under our experimental setup.
-Moreover, X-QL exhibited significant instability during the early stages of training in both robotic systems.
-Overall, these results demonstrate that SAC-BAR is better suited than the existing SOTA RL methods for the string-type artificial muscle robots targeted in this study.
+Moreover, X-QL exhibits significant instability during the early stages of training in both robotic systems.
+These results demonstrate that SAC-BAR is better suited than the existing SOTA methods for the string-type artificial muscle robots targeted in this study.
 VI-C Ablation Study in Simulation
-To evaluate the effectiveness of the proposed bootstrap and augmentation techniques, an ablation study was conducted on two robotic systems.
-Three variants of our proposed framework were tested as follows: (1) w/o A: This variant excludes the state vector augmentation method.
-(2) w/o B: This variant removes the bootstrap via PID method.
-(3) w/o AB: This baseline variant excludes both mechanisms.
-Each algorithm was trained and evaluated in simulation under randomized dynamics across four independent runs.
-Training results of our proposed SAC-BAR, with and without bootstrap and augmentation, are presented in Fig.
+To evaluate the effectiveness of the proposed bootstrap and augmentation techniques, an ablation study is conducted on two robotic systems.
+Five variants of our proposed framework are tested as follows: (1) Baseline: This variant employs only the SAC algorithm without any additional components.
+(2) w/ A: This variant incorporates the augmentation method.
+(3) w/ Ps: This variant perturbs the state vectors instead of target vectors.
+(4) w/ B: This variant integrates the bootstrap mechanism.
+(5) w/ Pt: This variant pre-trains the networks on data generated by the PID controller before exploration.
+Training results of these variants and our proposed SAC-BAR are presented in Fig.
 6(b).
 As illustrated in the figures, our proposed algorithm exhibits the highest level of data efficiency in both robotic systems.
-Compared to the SAC algorithm trained solely with dynamics randomization (referred to as “w/o AB”), the integration of bootstrap and augmentation enhances data efficiency and expedites the training process.
+Compared to the baseline SAC algorithm, the integration of bootstrap and augmentation enhances data efficiency and expedites the training process.
 In the robotic eye system, the baseline algorithm achieves an average reward of $-0.10$ after $2,000$ training episodes, whereas our approach achieves comparable performance in only approximately $750$ episodes, resulting in a $2.67$-fold increase in data efficiency compared to the original level.
 Similarly, in the robotic wrist system, the baseline algorithm achieves an average reward of $-1.61$ after $3,500$ episodes, while ours achieves comparable performance in only about $1,200$ episodes, leading to a $2.92$-fold increase in data efficiency.
-Notably, in the case of the more complex robotic wrist system, we find the demonstrations provided by the PID controller yield a more significant acceleration effect during the early stages of training.
-Meanwhile, the augmentation method demonstrates a favorable impact on enhancing the training speed in both systems and throughout the entire learning process of the algorithms.
+The performance of the variant perturbing state vectors is lower compared to the baseline.
+We attribute this to the perturbation of state variables, which likely hinders the agent’s ability to accurately learn the system dynamics.
+While pre-training on PID-generated data enhances training efficiency, we observe that overfitting may occur during the pre-training phase, leading to a significant drop in the average return.
+VI-D Influence of Hyperparameters
+Figure 7: (a) Comparative results for different scaling factors of the initial PID gains ($\beta=0.5,1.0,2.0,3.0$), where $K^{\prime}_{p}=\beta K_{p}$, $K^{\prime}_{i}=\beta K_{i}$, and $K^{\prime}_{d}=\beta K_{d}$.
+(b) Comparative results for different numbers of augmentation samples ($n=1,5,10,20$).
+The hyperparameters that significantly affect the performance of SAC-BAR include the initial PID gains and the number of augmentation samples.
+To evaluate their impact, we train the algorithm under different hyperparameter settings on the robotic wrist.
+Specifically, we compare the training results using different scaling factors for the initial PID gains ($\beta=0.5,1.0,2.0,3.0$) and varying numbers of augmentation samples ($n=1,5,10,20$), as shown in Fig.
+7.
+The results indicate that our method is generally robust to these hyperparameters.
+When the PID gains are within a reasonable range (not necessarily optimal), the differences in training results are minor.
+However, when the PID gains are extreme (e.g., $\beta=3$), the demonstrations it provides are of poor quality, thus affecting the algorithm’s training performance.
+Similarly, the performance of the model trained with $n=20$ drops significantly, as excessive augmentation introduces redundancy, causing the agent to overfit to the augmented data rather than focusing on the true dynamics of the environment.
+On the other hand, an extremely low $n$ reduces the diversity of the training data, slightly degrading performance.
+A moderate value of $n$ (e.g., $n=5$ or $n=10$) strikes a balance between data diversity and training stability.
 VII Real-World Experiments
-This section presents the real-world experimental results of the proposed SAC-BAR on two testing robotic platforms.
+This section presents the real-world experimental results of the proposed SAC-BAR.
 We first demonstrate the results of an ablation study conducted on the robotic wrist platform, which validate the effectiveness of the proposed dynamics randomization in enhancing control performance in real-world scenarios.
-Next, experimental results on the robotic eye are provided to show the generalizability of SAC-BAR across similar string-type artificial muscle-driven robotic systems.
-Additionally, we investigate the impact of the randomization range on control performance.
-VII-A Ablation Study on Robotic Wrist
+Next, experimental results on the robotic eye are provided to show the generalizability of SAC-BAR across similar robotic systems.
+Additionally, we investigate the impact of the randomization variance on control performance.
+VII-A Comparison Results on Robotic Wrist
 Figure 8: Comparison of experimental trajectories between the baseline and SAC-BAR.
 Target angles are set at ($\theta^{*}=2.5^{\circ},\phi^{*}=-7.5^{\circ}$).
-(a), (b) The system output angles ($\theta,\phi$) and corresponding actions $a_{t}$ of the baseline SAC algorithm.
+(a), (b) The system output angles ($\theta,\phi,e_{\text{ss}}$) and corresponding actions $a_{t}$ of the baseline SAC algorithm.
 (c), (d) Those of our proposed SAC-BAR algorithm.
@@
 In Fig.
-7(a), it is evident that without randomization of muscle dynamics, the vanilla-SAC struggles to achieve the desired orientation.
+8(a), it is evident that without randomization of muscle dynamics, the SAC agent struggles to achieve the desired orientation.
 While the agent may eventually reach the desired orientation at certain target angles, maintaining that target angle becomes a challenge, resulting in a considerable steady-state error.
 Conversely, as observed in Fig.
-7(c), the proposed SAC-BAR agent achieves the target state in less than $10$ s and consistently remains within its neighborhood thereafter.
+8(c), the proposed SAC-BAR agent achieves the target state in less than $10$ s and consistently remains within its neighborhood thereafter.
 As illustrated by the input action trajectories in Fig.
-7(d), this success is attributed to the agent learning to maximize the action value during the rising phase and promptly reduce it upon approaching the target angles.
+8(d), this success is attributed to the agent learning to maximize the action value during the rising phase and promptly reduce it upon approaching the target angles.
 Conversely, in Fig.
-7(b), there exist serious oscillations in the action trajectories of vanilla-SAC during the initial $10$ s, leading to a prolonged rise time and sometimes even failure to reach the target.
+8(b), there exist serious oscillations in the action trajectories of SAC during the initial $10$ s, leading to a prolonged rise time and sometimes even failure to reach the target.
 Figure 9: Experimental results of the steady-state error tests in the parallel robotic wrist under selected control algorithms.
@@
 (d) The box charts illustrating the medians, maxima, minima, and quartiles of the steady-state errors for the selected comparative control algorithms.
-In order to comprehensively evaluate the control performance of the algorithms across different target angles, we designed a steady-state error field test for the robotic wrist.
-We selected a series of target angles $(\theta^{*},\phi^{*})$ spaced at $2.5^{\circ}$ intervals within the defined range $\theta^{*},\phi^{*}\in[-10^{\circ},10^{\circ}]$.
-Each agent was set to run for $25$ s, sequentially targeting each angle pair.
+To comprehensively evaluate the control performance of the algorithms across different target angles, we design a steady-state error field test for the robotic wrist.
+We select a series of target angles $(\theta^{*},\phi^{*})$ spaced at $2.5^{\circ}$ intervals within the defined range $\theta^{*},\phi^{*}\in[-10^{\circ},10^{\circ}]$.
+Each agent is set to run for $25$ s, sequentially targeting each angle pair.
 The steady-state error is measured and averaged over the last $5$s window, determined by computing the Euclidean distance norm between the current and target angles, defined as $e_{\text{ss}}=\sqrt{(\theta-\theta^{*})^{2}+(\phi-\phi^{*})^{2}}$.
-The steady-state tracking errors for different targets of vanilla-SAC, SAC-R, and SAC-BAR are presented in Fig.
-8(a), (b), and (c), respectively.
+The steady-state errors for different targets of SAC, SAC-R, and SAC-BAR are presented in Fig.
+9(a), (b), and (c), respectively.
 The $\mathrm{x}$ and $\mathrm{y}$ axes therein represent $\theta^{*}$ and $\phi^{*}$, while the color bars represent the averaged steady-state errors.
 Following that, box charts are illustrated in Fig.
-8(d).
-These figures demonstrate that vanilla-SAC agent underperforms across nearly the entire defined range, whereas both SAC-R and SAC-BAR show significantly improved control performance.
-The steady-state errors averaged over the defined range, denoted as $\bar{e}_{\text{ss}}$, for three variants and the PID controller are presented in Table III.
-In real-world experiments, all three DRL-based controllers demonstrate superior performance compared to the traditional PID controller, reducing the tracking error by $11.6\%$, $35.9\%$, and $43.7\%$, respectively.
-Compared to the baseline SAC, SAC-R and SAC-BAR achieved reductions in the averaged tracking errors of $27.9\%$ and $36.3\%$, respectively.
-Specifically, our proposed SAC-BAR reaches the desired orientation at every target angle pair with an averaged error of $0.58^{\circ}$.
-Furthermore, as shown in Table III, the observed improvement in real-world performance can be attributed to overcoming the sim-to-real gap, as evidenced by the reduced disparity between simulated and actual results.
-TABLE III: The experimental (Real) and simulation (Sim) results of steady state error using different control methods
+9(d).
+These figures demonstrate that SAC agent underperforms across nearly the entire defined range, whereas both SAC-R and SAC-BAR show significantly improved control performance.
+These results indicate that the proposed randomization of muscle dynamics significantly enhances the real-world control performance.
+TABLE III: The simulation (Sim) and experimental (Real) results of steady state error using different control methods
 $\bar{e}_{\text{ss}}$ (deg) Control Method
-PID SAC SAC-R SAC-BAR
-Real $1.03\pm 0.49$ $0.91\pm 0.46$ $0.66\pm 0.39$ $0.58\pm 0.37$
-Sim $1.12\pm 0.47$ $0.46\pm 0.27$ $0.47\pm 0.33$ $0.40\pm 0.17$
-The above experimental results indicate that the proposed randomization of muscle dynamics significantly enhances the real-world control performance in the robotic systems of interest, while bootstrap and augmentation techniques contribute relatively less to the improvement.
-We believe that the enhancement observed in SAC-BAR compared to SAC-R primarily stems from the improvement in training efficiency.
-With the incorporation of bootstrap and augmentation methods, SAC-BAR achieves more comprehensive training and higher averaged rewards in simulation after 3,500 training episodes.
-We believe this improvement is potentially more significant in scenarios where exploration is more challenging.
+PID AC PPO SAC SAC-R SAC-BAR
+Sim 1.12±$\pm 0.47$ 0.80±$\pm 0.41$ 0.79±$\pm 0.32$ 0.46±$\pm 0.27$ 0.47±$\pm 0.33$ 0.40±$\pm 0.17$
+Real 1.03±$\pm 0.49$ 0.86±$\pm 0.67$ 1.24±$\pm 0.48$ 0.91±$\pm 0.46$ 0.66±$\pm 0.39$ 0.58±$\pm 0.37$
+To evaluate the real-world performance of our method against traditional control approaches, we implement the initial PID controller, an adaptive controller (AC), and the Proximal Policy Optimization (PPO) algorithm[42] for comparison.
+The steady-state errors averaged over the defined range, denoted as $\bar{e}_{\text{ss}}$, for these methods are tested in both simulation and real-world scenarios.
+The experimental results are summarized in Table III.
+In real-world experiments, our method achieves superior performance with a mean error of $0.58^{\circ}$, significantly lower than those of the other methods.
+Specifically, compared to PID, AC, PPO, and SAC, our method reduces the mean error by $43.7\%$, $32.6\%$, $53.2\%$, and $36.3\%$, respectively.
+While PPO and SAC are strong DRL baselines, they struggle with the sim-to-real gap, resulting in larger errors and less stable behaviors.
+The adaptive controller performs better due to its robustness to parameter uncertainties, but it requires extensive parameter tuning and still exhibits suboptimal performance at certain target angles.
+A comparison of SAC and SAC-R results demonstrates that the randomization of muscle dynamics effectively bridges the sim-to-real gap, leading to reduced disparity between simulated and actual results.
+Furthermore, SAC-BAR achieves more comprehensive training in simulation than SAC-R, resulting in a slight improvement in control performance.
 VII-B Learning Control of Robotic Eye
@@
 Target angles are set at $(\theta^{*}=-5^{\circ},\psi^{*}=7.5^{\circ})$ and $(\theta^{*}=10^{\circ},\psi^{*}=-5^{\circ})$, respectively.
-(a), (c) The system output angles $(\theta,\psi)$.
+(a), (c) The system output angles $(\theta,\psi,e_{\text{ss}})$.
 (b), (d) The corresponding actions $a_{t}$ generated by the trained actor.
-Figure 10: Experimental results on the steady-state tracking errors in robotic eye at eyery combination of $(\theta^{*},\psi^{*})$ target angles spaced at $2.5^{\circ}$ intervals for $(|\theta^{*}|,|\psi^{*}|)\leq 10^{\circ}$ where the color bar represents the Euclidean distance norm of steady-state error.
-To demonstrate the applicability of the proposed learning control framework across diverse robotic systems driven by artificial muscle strings, this subsection presents experimental results on an additional platform, the 2-DOF robotic eye.
-We focus on the foveation control task in the experiment.
+To demonstrate the applicability of the proposed learning control framework across diverse robotic systems driven by artificial muscle strings, this subsection presents experimental results on an additional platform.
 The robotic eye, initialized at $\theta=\psi=0^{\circ}$ at ambient room temperature, is configured to foveate at various target angles $(\theta^{*},\psi^{*})$.
@@
 These trajectories demonstrate that the measured angles rise to the target in approximately $5$ s, with negligible steady-state error.
-A similar field test is conducted with a series of target angles spaced at $2.5^{\circ}$ intervals within the defined range $|\theta^{*}|,|\psi^{*}|\leq 10^{\circ}$ to evaluate the overall control performance of the algorithm.
-The robotic eye is tested to foveate at each target in sequence.
+A similar field test is conducted to evaluate the overall control performance of the algorithm.
 The results, illustrated in Fig.
-10, reveal solid control performance, with steady-state errors consistently below $1^{\circ}$ for most target angles and an average of $0.56^{\circ}$.
-In our prior research[20], we employed a DRL-based controller using the Deep Deterministic Policy Gradient algorithm[43], which was trained in simulation for 50,000 episodes.
-Additionally, a reduced-order state observer was used to estimate the real-time temperature of SCP muscles, achieving satisfactory control performance.
+11, reveal solid control performance, with steady-state errors consistently below $1^{\circ}$ for most target angles and an average of $0.56^{\circ}$.
+Our prior research[17] employed a Deep Deterministic Policy Gradient-based controller[43], which was trained in simulation for $50,000$ episodes.
 In this paper, our proposed algorithm achieves comparable control performance with only $2,000$ training episodes, representing just $1/25$ of the original training time.
-Moreover, this is accomplished without the need to estimate real-time muscle temperature information under the condition of ambient temperature remaining relatively stable.
-VII-C Randomization Ranges on Control Performance
-TABLE IV: Comparison of control errors between models trained under different randomization ranges
-$\bar{e}_{\text{ss}}$ (deg) Randomization Range
-Small Original Large
-Real $0.63\pm 0.49$ $0.58\pm 0.37$ $0.76\pm 0.43$
-Sim $0.38\pm 0.17$ $0.40\pm 0.17$ $0.74\pm 0.41$
-In designing dynamics randomization, maintaining a balance between randomization level and model fidelity is crucial.
-Excessive randomization leads to an overly conservative controller, which may over-generalize and fail to leverage the more predictable aspects of system dynamics.
-Conversely, insufficient randomization causes the controller to overfit to a narrow set of conditions, thus limiting its performance when deployed in real-world systems.
-To investigate this trade-off, we conducted an experiment in which our control framework was trained with two additional randomization ranges—“small” and “large.” The small range was defined as half of the original randomization range, while the large range was doubled.
-We evaluated the performance of controllers trained under these varying randomization ranges in both simulated and real-world environments on the robotic wrist platform.
-The results, as summarized in Table IV, indicate that the control performance experiences a significant decline when randomization range is excessively large.
-Conversely, the smaller randomization range impairs the controller’s generalization capability from simulation to reality.
-Overall, a balanced range proves essential for ensuring robust performance.
+Figure 11: Experimental results on the steady-state errors in robotic eye at eyery combination of $(\theta^{*},\psi^{*})$ target angles spaced at $2.5^{\circ}$ intervals for $(|\theta^{*}|,|\psi^{*}|)\leq 10^{\circ}$.
+VII-C Different Randomization Variances
+TABLE IV: Comparison of control errors for models trained with different variances of the randomization distribution
+$\bar{e}_{\text{ss}}$ (deg) Standard Deviation Values
+$0.5\sigma_{0}$ $1.0\sigma_{0}$ $1.5\sigma_{0}$ $2.0\sigma_{0}$
+Sim 0.38±$\pm 0.17$ 0.40±$\pm 0.17$ 0.60±$\pm 0.35$ 0.74±$\pm 0.41$
+Real 0.63±$\pm 0.49$ 0.58±$\pm 0.37$ 0.64±$\pm 0.38$ 0.76±$\pm 0.43$
+To evaluate the influence of randomization variance on control performance, we conduct experiments using four standard deviation (SD) settings: $0.5\sigma_{0}$, $1.0\sigma_{0}$, $1.5\sigma_{0}$, and $2.0\sigma_{0}$, where $\sigma_{0}$ represents the baseline SD.
+Controllers trained under each randomization setting are evaluated in both simulated and real-world environments using the robotic wrist platform.
+As summarized in Table IV, the controller trained with a $2.0\sigma_{0}$ range exhibits a significant degradation in performance across both domains.
+In contrast, the controller trained with a $0.5\sigma_{0}$ range demonstrates limited generalization, failing to accommodate the variability inherent in real-world dynamics.
+While the selection of randomization variance significantly affects the resultant control performance, all controllers successfully converge during training.
+This indicates that the proposed method ensures reliable convergence across a reasonably bounded range of randomization settings.
+These results underscore the necessity of carefully tuning the randomization variance to balance generalization and performance without compromising convergence stability.
 VIII Conclusion
-This paper presents a learning-based control framework SAC-BAR designed for the string-type artificial muscle-driven robotic systems.
-To enhance training data efficiency and expedite convergence speed, we proposed two methods, namely bootstrap via PID and state-based data augmentation.
+This paper presents a learning control framework SAC-BAR designed for the string-type artificial muscle-driven robotic systems.
+To enhance training data efficiency, we proposed two methods, namely bootstrap via PID and state-based data augmentation.
 Additionally, we incorporated randomization of artificial muscle dynamic parameters during training to address the sim-to-real gap for enhancing real-world control performance.
-Extensive experiments and ablation studies on two systems—a robotic eye driven and a parallel robotic wrist—validated the effectiveness and generalizability of the proposed control framework.
+Extensive experiments and ablation studies on two robotic systems—a robotic eye and a parallel robotic wrist—validated the effectiveness and generalizability of the proposed control framework.
 In simulation, the designed bootstrap and augmentation methods reduced the required training episodes by approximately 60% compared to the baseline algorithm.
-Real-world experiments demonstrated that randomization of muscle dynamics improved the control performance of DRL-based controllers in targeted robotic systems, reducing the average error by 36.3%.
-Furthermore, our framework showed satisfactory performance on both robotic platforms, highlighting its potential applicability to similar artificial muscle robots.
+Real-world experiments demonstrated that randomization of muscle dynamics improved the control performance in targeted robotic systems, reducing the average error by 36.3%.
+Furthermore, our framework demonstrated satisfactory performance on both robotic platforms, highlighting its potential applicability to other robotic systems with actuators characterized by high nonlinearity and parameter uncertainties.
 Despite the strong performance of the proposed SAC-BAR, several limitations exist.
-The configuration of the reward function has a significant impact on experimental outcomes, often requiring empirical tuning to ensure optimal controller performance.
-The dynamic models of the artificial muscles and robotic wrist are simplified and do not fully represent their true dynamics.
-In future work, we plan to explore automated methods for reward function tuning, potentially incorporating adaptive or self-tuning mechanisms that respond dynamically to performance metrics.
-To better account for real-world variability, we will extend the randomization of muscle dynamics by incorporating high-fidelity models that capture nonlinearity and time-dependent factors, using data-driven or hybrid modeling approaches.
+The configuration of the reward function relies on empirical tuning, and this paper does not provide a formal theoretical convergence analysis under muscle dynamic randomization.
+Additionally, the application scenarios explored in this work are relatively limited, and the robotic testing platforms utilized share similar characteristics.
+In future work, we aim to explore automated methods for reward function tuning and develop theoretical convergence analysis frameworks that consider bounded randomization ranges and non-stationary dynamics.
+To further evaluate the practicality of the proposed method, we plan to extend its application to more complex systems and diverse scenarios.